### v1

# A Langevin sampling algorithm inspired by the Adam optimizer

Abstract
We present a framework for adaptive-stepsize MCMC sampling based on time-rescaled Langevin dynamics, in which the stepsize variation is dynamically driven by an additional degree of freedom. Our approach augments the phase space by an additional variable which in turn defines a time reparameterization. The use of an auxiliary relaxation equation allows accumulation of a moving average of a local monitor function and provides for precise control of the timestep while circumventing the need to modify the drift term in the physical system. Our algorithm is straightforward to implement and can be readily combined with any off-the-peg fixed-stepsize Langevin integrator. As a particular example, we consider control of the stepsize by monitoring the norm of the log-posterior gradient, which takes inspiration from the Adam optimizer, the stepsize being automatically reduced in regions of steep change of the log posterior and increased on plateaus, improving numerical stability and convergence speed. As in Adam, the stepsize variation depends on the recent history of the gradient norm, which enhances stability and improves accuracy compared to more immediate control approaches. We demonstrate the potential benefit of this method–both in accuracy and in stability–in numerical experiments including Neal’s funnel and a Bayesian neural network for classification of MNIST data.
Keywords Sampling methods, computational statistics, Adam, Langevin dynamics, adaptive or variable stepsize, Bayesian sampling, neural network
1 Introduction
Monte Carlo sampling schemes are ubiquitous in modern-day science, engineering, and finance. They are used to quantify risk and uncertainty, parameterize statistical models, calculate thermodynamic quantities in physical models, and explore protein configurational states. The aim in sampling is to generate independent and identically distributed (i.i.d.) realizations $x_{i}$ of a random variable $X\in\mathbb{R}^{d}$ distributed according to a given probability law, which we assume is defined in terms of a positive, smooth density $\pi$. Such samples can then be used to estimate probabilities of certain events or expectations of functions, to calculate uncertainties, assess or compare models, or to explore optimality of parameterizations in machine learning applications. One of the most powerful and widely used categories of sampling methods are Markov Chain Monte Carlo schemes (MCMC [3, 13]), which generate the samples using Markov Chains that are ergodic with respect to the target probability measure. The original MCMC scheme is the Metropolis-Hastings (MH) algorithm [66, 40] which uses random proposals in conjunction with an accept-reject procedure (the MH-criterion) to generate the Markov Chain. The MH framework is very general and allows many alternatives for proposal generation. The use of Metropolis correction may, however, add substantial computational burden and the rejection steps can slow convergence to the target distribution.
In this article we consider Langevin dynamics-based methods. We assume the target probability density can be defined in terms of a $C^{2}(\mathbb{R}^{d})$ energy function $U:\mathbb{R}^{d}\to\mathbb{R}$ which grows sufficiently rapidly as $x\rightarrow\infty$ so that
$\pi(x)=Z^{-1}\exp(-\beta U(x))$
[…]
${\rm d}x=-\nabla U(x){\rm d}t+\sqrt{2\beta^{-1}}{\rm d}W_{t}$ (1)
is used to generate the paths, where $(W_{t})_{t\geq 0}$ is a $d$-dimensional standard Brownian motion. In practice, it is found that introducing a momentum vector $p$ can enhance the efficiency of the sampling process and the “underdamped” Langevin dynamics system is often used instead:
${\rm d}x$ $=$ $p{\rm d}t,$ (2)
${\rm d}p$ $=$ $-\nabla U(x){\rm d}t-\gamma p{\rm d}t+\sqrt{2\gamma\beta^{-1}}{\rm d}W_{t}.$ (3)
Convergence analysis of (1) and (2)-(3) is well understood in the continuous setting [75], but these continuous systems need to be replaced by discrete processes in practical applications. For this purpose a numerical discretization is introduced. For example, (1) can be discretized using the Euler-Maruyama method:
$x_{n+1}=x_{n}-\Delta t\nabla U(x_{n})+\sqrt{2\Delta t\beta^{-1}}\xi_{n+1},$
where $\Delta t>0$ is the stepsize and $\xi_{n+1}\sim\mathcal{N}(0,I_{d})$. Similar discretizations may be introduced to solve (2)-(3). For example one method maps $(x_{n},p_{n})$ to $(x_{n+1},p_{n+1})$ by
$p_{n+1}$ $=$ $cp_{n}-\Delta t\nabla U(x_{n})+\sqrt{(1-c^{2})\beta^{-1}}\xi_{n+1},$
$x_{n+1}$ $=$ $x_{n}+\Delta tp_{n},$
where $c=\exp(-\Delta t\gamma)$ and $\xi_{n+1}\sim\mathcal{N}(0,I_{d})$; this is referred to as the “OBA” method using the naming convention first introduced in [55]. While the momenta must be carried forward during computations, only the sequence of $x$ variables needs to be stored.
By eliminating the stop-and-go aspect of MH methods (i.e., rejected steps), SDE schemes promise faster convergence, but they have some important drawbacks. First, a numerical method such as those mentioned above introduces bias with respect to the target distribution. Effectively we sample a perturbed distribution with density $\hat{\pi}^{\Delta t}$ replacing $\pi$. The bias can be controlled by choosing a sufficiently small discretization stepsize $\Delta t$, where the quality of the numerical integrator governs the size of the bias and its stepsize-dependent scaling [57]. Of course, the stepsize cannot be reduced arbitrarily, as smaller steps lead to more strongly correlated successive states. This, in turn, requires longer trajectories to achieve the same level of exploration, increasing the computational cost.
The second well-known drawback of LD methods is that such methods tend to suffer from numerical instabilities in cases where $\nabla U$ has a large Lipschitz constant, forcing the stepsize to be reduced. The stepsize restriction will be governed by the largest curvature; in the case of multimodal target distributions each basin may have a different Hessian eigenstructure and thus introduce different stability constraints. As the sampling path ventures from the vicinity of one minimum to the vicinity of another, the stepsize restrictions change. In common practice a single fixed stepsize is used which must be chosen small enough to mitigate these issues. This comes at the cost of requiring a larger number of integrator steps to generate a new (sufficiently decorrelated) sample, thus increasing the computational cost in comparison to what would seem intuitive.
In high-dimensional sampling applications, individual integrator steps may be extremely expensive due to the costly evaluation of $\nabla U$. Nowhere is this more apparent than in machine learning, where a reduction of computing cost per iteration is typically achieved by replacing the evaluation of the true gradient $\nabla U$ with a cheaper stochastic approximation (usually realized by data subsampling, see e.g., [82, 100, 20, 30, 89]). This, however, adds additional perturbations to the dynamics which can again be controlled by decreasing the stepsize (see e.g., [87] for a simple model of how gradient noise relates to stepsize). Thus, substantial effort has been made to design novel LD-based MCMC methods with improved accuracy, stability, or efficiency, see, for example [15, 65, 16, 48, 55, 74, 100, 20, 30, 23, 88, 39]. It should be mentioned that there are also dynamics-based MCMC schemes that come with an MH-correction step, such as Hamiltonian Monte Carlo (HMC) and relatives [32, 13, 42, 12, 38, 83, 81]. However, due to the expensive evaluation of $U$ and potentially low acceptance rates, they are often deemed too inefficient for large scale applications [100, 20, 6]. Efforts to address the issue [103, 104] may sacrifice stability or robustness compared to standard procedures and have not been widely adopted. This article proposes an unadjusted LD-based sampling scheme and is thus in line with the philosophy of foregoing the MH-criterion.
One approach to tackle some of the mentioned issues comes in the form of adaptive stepsize methods, which are widely used in integration of deterministic systems, e.g., $\frac{\mathop{}\!\text{d}}{dt}x=\phi(x)$. Variable stepsize procedures try to estimate the local discretization error being introduced at each step and adjust the stepsize down or up to maintain a certain prescribed local error tolerance (see, e.g., Chapter 3 of [54]). The error estimate may be based on finite difference approximation, or extrapolation, or the use of specialized ‘embedded’ integration schemes (e.g., Runge-Kutta Fehlberg methods [36]). These techniques have also been adapted for weak approximation (sampling) using stochastic differential equations [84, 97, 94]. Implementations of this approach are in widespread use in modern simulation software. The molecular dynamics package OpenMM [35] uses a variable stepsize scheme in which local errors are estimated on-the-fly via simple Euler discretization (or Euler-Maruyama discretization for their LD integrator; see [35]). Additionally, there has been recent developments of adaptive stepsize methods, which allow for high strong order approximation in [47] and [37]. We also mention the development of recent methodology for local stepsize adjustment for HMC based on Gibbs self tuning in [9] and local adaptation of other parameters in HMC [10].
An alternative approach to variable stepsize changes the stepsize through Sundman time transformation:
$\frac{\mathop{}\!\text{d}}{\mathop{}\!\text{d}\tau}x=R(x)\phi(x),\quad\frac{\mathop{}\!\text{d}t}{\mathop{}\!\text{d}\tau}=R(x),$ (4)
with a scalar-valued function $R:\mathbb{R}^{d}\to\mathbb{R}$, which is uniformly bounded such that $0<m<R(x)<M<\infty$ for all $x\in\mathbb{R}^{d}$. These types of transformations are commonly used in classical mechanics (see the recent work [17] and the references therein, as well as their illustrative examples). The Sundman transform can be used to turn fixed-stepsize numerical integrators into adaptive stepsize schemes improving integration stability and efficiency. Specifically, one may discretize the equation (4) using a fixed step $\Delta\tau$ in the ‘fictive’ time variable $\tau$, then interpret this as equivalent to variable steps in ‘real’ time according to (at timestep $n$)
$\Delta t_{n}\approx R(x_{n})\Delta\tau.$
Recently, the idea was applied to Langevin dynamics in [61] and more general Markov processes [5]. In [61], a suggested transform kernel was given by $R(x)=\tilde{R}\left(\|\nabla U(x)\|^{-1}\right)$, with a boundedness-ensuring function $\tilde{R}:\mathbb{R}_{+}\to\left[m,M\right]$ with $\lim_{u\to\infty}\tilde{R}(u)=m<M=\lim_{u\to 0}\tilde{R}(u)$. As in the ODE case, the corresponding numerical schemes accomplish enhanced stability and efficiency on the one- and two-dimensional test examples considered. The framework we present in this article builds on this idea, but we introduce an alternative mechanism for stepsize adaptation. Rather than using the current state of the variable of interest $x$ to adjust the stepsize, we introduce an auxiliary variable $\zeta\in\mathbb{R}$, evolving via a suitably chosen dynamical equation
$\frac{\mathop{}\!\text{d}}{\mathop{}\!\text{d}\tau}\zeta=f(x,p,\zeta).$ (5)
We focus here on the following natural choice:
$f(x,p,\zeta)=-\alpha\zeta+g(x,p),$
with parameter $\alpha>0$ and a monitor function $g$. This type of dynamics effectively computes a moving average of $g(x,p)$ over the recent history, with the driving function $g$ determining on what basis the stepsize is to be changed. We then apply a Sundman transformation which is expressed as a function of $\zeta$
${\rm d}t=\psi(\zeta){\rm d}\tau.$
As the time-rescaling alters the rate at which samples are acquired, these samples cannot be used directly for computing Gibbs-Boltzmann averages, but as we shall see it is straightforward to reweight the data in order to calculate any such quantities. Employing a bounded Sundman transformation $\psi$ controls the stability of the reweighting process.
For both the introduction of the general dynamics (5) and the choice of $g$, we draw inspiration from optimization, where numerical challenges arise that mirror those associated to MCMC samplers. A classic, general purpose optimization method is the stochastic gradient descent (SGD) method of Robins and Monro [82], and it remains a popular choice due to its simplicity and robustness in practice. In recent years, there has been extensive work on adaptive-stepsize variants of SGD [33, 96, 102, 53, 31, 79, 21]. Most of these schemes are modifications or extensions of the methods AdaGrad [33, 64] or RMSProp [96]. By far the most widely known and used of these is the Adam optimizer (short for adaptive moment estimation [53]), which uses estimates of mean and variance of the gradients (computed as moving averages) to adjust the stepsize, reducing it in relation to the steepness of the landscape. Adam is known to improve efficiency in certain cases [53], often reaching the minimum in many fewer steps than SGD, and is popular for model training in various settings such as natural language processing [98, 29, 14], Bayesian neural networks [44, 62], and computer vision [77, 45, 52].
By combining the methodology of time-transformed SDEs with Adam’s approach to adjust the stepsize based on a moving average over recent history, we derive a flexible framework for adaptive-stepsize sampling, which we call “SamAdams” (sampling with adaptive moderated stepsize). The procedure can be combined with state-of-the art integrators for Langevin dynamics. In particular, it is possible to transform the splitting integrators mentioned above into adaptive algorithms. Adaptivity improves numerical stability compared to relying on a constant stepsize, as demonstrated in Fig. 1 where a fixed stepsize Langevin integrator is compared with its adaptive counterpart.
Figure 1: Sampling trajectories of a constant-stepsize integrator (BAOAB) and our adaptive-stepsize scheme (SamAdams) on a star-shaped landscape $U(x,y)=x^{2}+1000x^{2}y^{2}+y^{2}$. Left: Potential $U(x,y)$ with trajectories. BAOAB was run at the mean stepsize used by SamAdams (obtained by averaging over all iterations). Right: Stepsize values $\Delta t$ used by SamAdams are binned by distance to the origin $r=\sqrt{x^{2}+y^{2}}$ together with the mean stepsize (blue dashed line) and the maximum stable stepsize for BAOAB (red dashed line). SamAdams uses a small stepsize only at the outer points of the stable domain.
While the fixed stepsize method becomes unstable in the tips of the star-like potential, SamAdams remains stable by automatically reducing its stepsize in those regions. As evident by the $\Delta t$ histograms, the smallest $\Delta t$ values are adopted in the tips of the star. There, they are similar in size to the stability threshold of BAOAB, which neatly confirms the understanding that the stability of a constant-stepsize scheme is determined by those landscape regions of largest steepness and curvature where the forces and force fluctuations are largest. Since most of the probability mass is located close to the origin where the experienced forces are small, the adaptive scheme is able to use larger stepsizes during most of the simulation, enabling a larger mean stepsize $\langle\Delta t\rangle$, with consequent increase in computational efficiency. The ability to use larger stepsizes while maintaining sampling quality would greatly benefit those sampling applications where the sampling error is dominated by a lack of exploration of the loss landscape $U$ (rather than other sources of error such as discretization bias or model specification). Prominent examples are large-scale Bayesian neural networks [71, 70, 49, 101, 46] or molecular dynamics simulations [86, 41, 56, 25, 35].
For the purposes of disambiguation we mention that Adam and variants further improve optimizer efficiency by independently rescaling the coordinates of the system through the mechanism of individual timesteps. We don’t address this important aspect here but instead focus on the use of a moving average to stabilize the timestep selection in a pure Langevin dynamics framework. Some sampling methods that incorporate anisotropic coordinate transformation in order to enhance performance are RMHMC [38], the ensemble quasi-Newton method [58], and the recently proposed AdamMCMC method [7].
The rest of this article is structured as follows. Section 2 discusses the Sundman-transformed SDE our new framework is built on and a reweighting scheme for physical observables from trajectories that evolve in rescaled time. The following section is addressed to numerical discretization of the equations of motion. In Section 4 we discuss designing a Sundman kernel so that the transformed SDE adopts a stepsize adaptation resembling the device in Adam. Finally, Section 5 contains our numerical results.
2 Sundman-transformed SDEs and Averaging
The notion of rescaling of time is a familiar one in studies of gravitational $N$-body problems, where it is used in analytical as well as numerical treatments. Let an autonomous ordinary differential equation be given of the form
$\frac{\mathop{}\!\text{d}x_{t}}{\mathop{}\!\text{d}t}=\phi(x_{t}).$ (6)
We use the notation $x_{t}:=x(t)$ to compactly indicate the independent variable. In particular, the classical Sundman transformation[93] replaces $t$ by a new variables $\tau$ which is defined by an ordinary differential equation of the form
$\frac{\mathop{}\!\text{d}t}{\mathop{}\!\text{d}\tau}=\psi(x_{\tau}),$
with $x_{\tau}:=x_{t(\tau)}$, so that (6) becomes, using chain rule,
$\frac{\mathop{}\!\text{d}x_{\tau}}{\mathop{}\!\text{d}\tau}=$ $\frac{\mathop{}\!\text{d}x_{t}}{\mathop{}\!\text{d}t}\frac{\mathop{}\!\text{d}t}{\mathop{}\!\text{d}\tau}=\psi(x_{\tau})\phi(x_{\tau}).$
This type of time-rescaling can be used to facilitate numerical integration of ODEs. As explained in [92, 43], if the Sundman transformation is suitably chosen to normalize the system or at least lessen the variation in the magnitude of $\phi$ and its derivatives, the rescaled system can often be integrated using fixed stepsize with the result that errors or instabilities associated with directly integrating (6) are eliminated or reduced.
Time-rescaling changes the effective frequencies of a system with oscillatory components. For example if we introduce a simple constant Sundman transformation ${\rm d}t/{\rm d}\tau=a$ into a 1D harmonic oscillator with frequency $\omega$, $\mathop{}\!\text{d}q_{t}/\mathop{}\!\text{d}t=p_{t}$; $\mathop{}\!\text{d}p_{t}/\mathop{}\!\text{d}t=-\omega^{2}q_{t}$, we obtain the system $\mathop{}\!\text{d}q_{\tau}/\mathop{}\!\text{d}\tau=ap_{\tau}$; $\mathop{}\!\text{d}p_{\tau}/\mathop{}\!\text{d}\tau=-a\omega^{2}q_{\tau}$. This new system has frequency $a\omega$. The same effect can be obtained by rescaling only the momentum equation or only the position equation. Since the stable timestep for integration of oscillatory dynamics typically depends on the fast frequencies we can potentially improve stability efficiently by adjusting the time-rescaling dynamically. In nonlinear systems and systems with multiple oscillatory modes, a configuration-dependent Sundman transformation can be used to modify the dynamics so that all components are propagated with an effective smaller stepsize in regions of high oscillation. For example, the Adaptive Verlet method [43] uses just such a time-rescaling, treating the local time-rescaling factor as an additional dependent variable of the system and evolving this in various ways to enhance numerical performance.
A simpler form of adaptivity that, in our experience, often works well, is to use the Sundman transformation to adjust the stepsize at the beginning of each step using $\Delta t_{n}:=\Delta\tau R(x_{n})$. $\Delta\tau$ represents a fixed stepsize for the rescaled system and $R(x)$ indicates how the step should be adjusted depending on current configuration $x$. As in [61], by applying the Sundman transform to an SDE and then discretizing with fixed stepsize, we can obtain an adaptive stepsize sampling method in the original time variable.
Using the Sundman transform $\mathop{}\!\text{d}t/\mathop{}\!\text{d}\tau=R(x_{\tau})$, we can write down a time-rescaled version of (2)-(3):
$\mathop{}\!\text{d}x_{\tau}$ $=R(x_{\tau})p_{\tau}\mathop{}\!\text{d}\tau,$ (7)
$\mathop{}\!\text{d}p_{\tau}$ $=-R(x_{\tau})\nabla U(x_{\tau})\mathop{}\!\text{d}\tau-\gamma R(x_{\tau})p_{\tau}\mathop{}\!\text{d}\tau+\sqrt{\frac{2\gamma R(x_{\tau})}{\beta}}\mathop{}\!\text{d}W_{\tau}.$ (8)
In [61], it was shown that the canonical distribution $\rho_{\beta}$ is no longer invariant under the process, so that work introduced an additional drift correction term to recover useful samples.
Here, we propose a different way of introducing the timestep adaptation. Rather than letting the transform function directly depend on the configuration $x$, we introduce an artificial dynamical control variable $\zeta\in\mathbb{R}$, and a Sundman rescaling function $\psi$ such that the time adaption is governed by $\mathop{}\!\text{d}t/\mathop{}\!\text{d}\tau=\psi(\zeta_{\tau})$ with $\mathop{}\!\text{d}\zeta_{\tau}=f(x_{\tau},p_{\tau},\zeta_{\tau})\mathop{}\!\text{d}\tau$ for some scalar-valued function $f$ on the augmented phase space variables $(x,p,\zeta)$.
It makes intuitive sense to choose the function $f$ as the sum of dissipative and driving terms, i.e.,
$f(x,p,\zeta)=-\alpha\zeta+g(x,p),$ (9)
with hyperparameter $\alpha>0$ representing the attack rate of a relaxation process. The monitor function $g(x,p)$ can be any positive smooth function and ultimately decides on what basis a small time increment $\mathop{}\!\text{d}t$ is varied. The rescaling function $\psi$, over a crucial interval, mimics a reciprocal power of $\zeta$. If the monitor function takes on larger values, $\zeta$ will tend to increase and the time interval $\mathop{}\!\text{d}t$ will shrink (implying smaller stepsizes during numerical integration). If the monitor function is small (meaning the solution is locally smooth and easy to integrate), the rescaling will lead to an increase in stepsize. This corresponds to the paradigm to take as large a stepsize as possible but as small as necessary. The dynamics (9) effectively computes a moving average of the driving function $g$ over the recent history, exponentially weighted with rate $\alpha$ (see the discussion in Appendix C). Making the stepsize adaptation depend on the recent history of a driving function is one reason why the Adam optimizer became so successful.
Since one can employ many different monitor functions $g(x,p)$ and Sundman transform kernels $\psi(\zeta)$, we obtain a flexible framework for adaptive stepsize sampling, given by [^1] A version of this system for generic SDEs is given in the Supplementary Material.
$\mathop{}\!\text{d}x_{\tau}$ $=\psi(\zeta_{\tau})p_{\tau}\mathop{}\!\text{d}\tau,$ (10)
$\mathop{}\!\text{d}p_{\tau}$ $=-\psi(\zeta_{\tau})\nabla U(x_{\tau})\mathop{}\!\text{d}\tau-\gamma\psi(\zeta_{\tau})p_{\tau}\mathop{}\!\text{d}\tau+\sqrt{2\gamma\beta^{-1}\psi(\zeta_{\tau})}\mathop{}\!\text{d}W_{\tau},$ (11)
$\mathop{}\!\text{d}\zeta_{\tau}$ $=-\alpha\zeta_{\tau}\mathop{}\!\text{d}\tau+g(x_{\tau},p_{\tau})\mathop{}\!\text{d}\tau,$ (12)
$\mathop{}\!\text{d}t$ $=\psi(\zeta_{\tau})\mathop{}\!\text{d}\tau.$ (13)
In Sec. 4 we show that we can pick the Sundman transform $\psi$ and driving function $g$ to obtain a sampler that resembles the Adam optimizer, although it is important to emphasize that the framework is very general. We next address the ergodicity of process (10)-(13), how to obtain canonical averages from it, and how to simulate it in practice.
2.1 Ergodic Averages
Due to the time-rescaling and the additional $\zeta$-dynamics, the invariant measure of (10)-(13) no longer coincides with the invariant measure of underdamped Langevin dynamics (2)-(3), i.e., the canonical measure $\pi_{\beta}$. However, under mild assumptions on $U$ and $\psi(u)$, we can correct this error by reweighting the samples with the corresponding values of the transform kernel $\psi(\zeta_{\tau})$. To see this, assume the dynamics (10)-(13) is ergodic with invariant measure $\Pi_{\tau}$. Assume we have samples $(x_{\tau_{i}},p_{\tau_{i}},\zeta_{\tau_{i}})$ from the solution of (10)-(13) at times $\tau_{i}:=i\Delta\tau$, $i\in\mathbb{N}$, for some stepsize $\Delta\tau>0$. It follows that under mild assumptions on an observable $\phi:\mathbb{R}^{2d}\to\mathbb{R}$,
$\lim_{N\to\infty}\frac{\sum^{N}_{i=1}\phi(x_{\tau_{i}},p_{\tau_{i}})\psi(\zeta_{\tau_{i}})}{\sum^{N}_{i=1}\psi(\zeta_{\tau_{i}})}$ $=\lim_{N\to\infty}\frac{\frac{1}{N}\sum^{N}_{i=1}\phi(x_{\tau_{i}},p_{\tau_{i}})\psi(\zeta_{\tau_{i}})}{\frac{1}{N}\sum^{N}_{i=1}\psi(\zeta_{\tau_{i}})}$
$=\frac{\mathbb{E}_{\Pi_{\tau}}(\phi\psi)}{\mathbb{E}_{\Pi_{\tau}}(\psi)}$
[…]
$=\mathbb{E}_{\pi_{\beta}}(\phi),$
where the second and third lines follow from ergodicity and the fourth line from identifying $\psi\mathop{}\!\text{d}\tau=\mathop{}\!\text{d}t$. We remark that the process (10)-(13) can be shown to be ergodic under minimal assumptions as in [85] which treats a more general case. Similar techniques can be found in [63, 61, 76]. In particular, under sufficient smoothness, the assumption that the time-rescaling $\psi$ is uniformly bounded from below and above and the force is convex outside a ball, one can show ergodicity of the time-rescaled process (10)-(13) by considering the family of Lyapunov functions considered in [63]. Resulting in an ergodicity result with a convergence rate which depends on the uniform lower bound on $\psi$. More sophisticated techniques would be required to show an improved convergence rate of the Adam sampler, akin to the adaptive stepsize results available in the optimization literature (see, for example, [27] for Adam or Adagrad).
We have thus shown how to obtain canonical averages as time averages over a single trajectory. In practice, when employing constant-stepsize MCMC to sample $\pi_{\beta}$, rather than obtaining averages through a time average along a single long trajectory, one often draws multiple trajectories in parallel (allowing for parallel computation) and approximates
$\mathbb{E}_{\pi_{\beta}}(\phi(x_{t},p_{t}))\approx\frac{1}{N}\sum_{i=1}^{N}\phi(x^{i}_{t},p^{i}_{t}),$ (14)
for $N$ independent trajectories and $t$ large enough to have $\text{Law}(x_{t},p_{t})\approx\pi_{\beta}$. The superscript refers to the trajectory index. The same is possible for the time-rescaled dynamics (10)-(13). For sufficiently large $\tau(t)$, we have that $\text{Law}(x_{\tau},p_{\tau},\zeta_{\tau})\approx\Pi_{\tau}$. In this case, we have for $N$ independent trajectories, $N$ large enough,
$\frac{\sum_{i=1}^{N}\phi(x^{i}_{\tau},p^{i}_{\tau})\psi(\zeta^{i}_{\tau})}{\sum_{i=1}^{N}\psi(\zeta^{i}_{\tau})}\approx\frac{\mathbb{E}_{\Pi_{\tau}}(\phi R)}{\mathbb{E}_{\Pi_{\tau}}(R)}=\mathbb{E}_{\pi}(\phi).$ (15)
Thus, when simulating the time-rescaled dynamics, one can use both time- and trajectory- averages to approximate canonical averages just like in constant-stepsize MCMC. Obtaining canonical averages can thus be illustrated by the diagram below.
Figure 2: SamAdams sampling procedure.
The subscript $i$ in the upper right-hand corner of the diagram can either denote different points in time along a single trajectory (time-average) or a trajectory index when averaging over samples of different trajectories taken at the same time (trajectory-average).
3 Numerical Integration
In order to simulate the time rescaled dynamics (10) to (13), we need to discretize the continuous-time process with a suitable numerical integrator. For convenience we adopt the framework of [91, 16, 11, 55] in which a symplectic splitting of the Hamiltonian part of the underdamped Langevin system is composed with a map that exactly preserves the momentum distribution. A frequent choice for the stochastic part is
$\Phi_{\Delta t}^{{\rm O}}(x,p)=\bigg{(}x,\ \exp(-\gamma\Delta t)p+\sqrt{(1-\exp(-2\gamma\Delta t))\beta^{-1}}\xi\bigg{)},$
where $\xi$ is a random vector with each component independently drawn from ${\mathcal{N}}(0,1)$. Since $\Phi_{\Delta t}^{{\rm O}}$ preserves the momentum distribution regardless of $\Delta t$, it is reasonable to introduce this in a way which parallels our previous derivation. Resolving the Hamiltonian part can be done by splitting the dynamics into a drift at constant momentum $\mathop{}\!\text{d}x_{t}/\mathop{}\!\text{d}t=p_{t}$; $\mathop{}\!\text{d}p_{t}/\mathop{}\!\text{d}t=0$ and a momentum kick $\mathop{}\!\text{d}x_{t}/\mathop{}\!\text{d}t=0$; $\mathop{}\!\text{d}p_{t}/\mathop{}\!\text{d}t=-\nabla U(x_{t})$ using the two propagators
$\Phi_{\Delta t}^{{\rm A}}(x,p)$ $=(x+\Delta tp,\ p),$
$\Phi_{\Delta t}^{{\rm B}}(x,p)$ $=(x,\ p-\Delta t\nabla U(x)).$
Different integrators for Langevin dynamics can be obtained by composing the A-, B-, and O-maps in different ways, e.g., giving symmetric OBABO [16] and BAOAB[55] schemes; for example $\Phi^{\text{OBABO}}_{\Delta t}:=\Phi_{\Delta t/2}^{{\rm O}}\circ\Phi_{\Delta t/2}^{{\rm B}}\circ\Phi_{\Delta t}^{{\rm A}}\circ\Phi_{\Delta t/2}^{{\rm B}}\circ\Phi_{\Delta t/2}^{{\rm O}}$. In [55, 57] these integrators have been shown to provide second order approximation of averages with respect to the Gibbs-Boltzman distribution. Because of the special form of these composition methods, only a single evaluation of the expensive gradient term is needed at each step, improving efficiency. Alternatives such as BABO (half step of B, whole step of A, half step of B, whole step of O) sacrifice the symmetry but maintain the same second order accuracy with respect to the invariant measure and similarly require only a single gradient per step. Another good prospective Langevin scheme is the symmetric UBU discretization [90].
What remains is to incorporate the stepsize adaptation described by (12) and (13). In a splitting approach, the auxiliary variable $\zeta$ evolves using (12) with fixed $x$ and $p$, thus the solution is just
$\zeta(\tau)=e^{-\alpha\tau}\zeta(0)+g(x,p)\int_{0}^{\tau}e^{-\alpha(\tau-s)}{\rm d}s.$ (16)
Hence we may introduce an additional map
$\Phi^{\rm Z}(x,p,\zeta)=\left(x,\ p,\ \exp(-\Delta\tau\alpha)\zeta+\frac{1}{\alpha}\left(1-e^{-\Delta\tau\alpha}\right)g(x,p)\right).$ (17)
Once $\zeta$ has been calculated, the new stepsize $\Delta t$ is updated via $\Delta t:=\psi(\zeta)\Delta\tau$. In Sec. 4 we introduce two suitable choices for the transform kernel $\psi$.
With $x,p$ held fixed and thus $g(x,p)\equiv g$, we may rewrite the update for $\zeta$ as follows,
$\zeta_{n+1}=\hat{\Phi}^{\rm Z}(x,p,\zeta_{n})\equiv\rho\zeta_{n}+\alpha^{-1}(1-\rho)g(x,p),$
where $\rho=\exp(-\alpha\Delta\tau)$, which is fixed along a sampling path.
The choice $\rho\approx 1$ maintains a strong dependence on the stepsize history, whereas $\rho\approx 0$ represents a rapid damping. We also define
$\hat{\Phi}^{Z}_{a}(x,p,\zeta)=\rho^{a}\zeta+\alpha^{-1}(1-\rho^{a})g(x,p),$
to allow for taking partial steps of the $\zeta$ flow.
3.1 The Algorithm
A procedure for implementation of a symmetric variant of SamAdams is given in Alg. 1.
Algorithm 1 SamAdams
Given: $\hat{\Phi}_{\Delta t}$ a (fixed-stepsize) Langevin integrator, ${\psi}$ a suitable Sundman transformation.
Given: parameters $n_{\rm max}$, $\Delta\tau$, and $n_{\text{meas}}$.
Given: initial conditions $x_{0}$, $p_{0}$, $\zeta_{0}$. Set $\mu_{0}={\psi}(\zeta_{0})$.
for $n=0:n_{{\rm max}}$ do
     if $n\ \text{mod}\ n_{\text{meas}}==0$ then
         Collect sample $(x_{n},p_{n},\mu_{n})$. $\triangleright$ $\mu_{n}$ needed for reweighting.
     end if
     $\zeta_{n+\frac{1}{2}}:=\hat{\Phi}^{Z}_{1/2}(x_{n},p_{n},\zeta_{n})$. $\triangleright$ Evolve $\zeta$.
     $\Delta t_{n+1}:={\psi}(\zeta_{n+\frac{1}{2}})\Delta\tau$. $\triangleright$ Modify stepsize.
     $({x}_{n+1},{p}_{n+1}):=\hat{\Phi}_{\Delta t_{n+1}}(x_{n},p_{n})$. $\triangleright$ Langevin dynamics step with stepsize $\Delta t_{n+1}$.
     $\zeta_{n+1}:=\hat{\Phi}^{Z}_{1/2}(x_{n+1},p_{n+1},\zeta_{n+\frac{1}{2}})$. $\triangleright$ Evolve $\zeta$.
     $\mu_{n+1}:={\psi}(\zeta_{n+1})$. $\triangleright$ Calculate weight.
end for
We can name integrators from the SamAdams family simply by introducing letters Z corresponding to updates of the $\zeta$ dynamics. Palindromic letter sequences like ZBAOABZ indicate a symmetric introduction of $\zeta$ updates around a fixed stepsize method. Compared to the alternative ZBAOAB, the symmetric inclusion of the two Z half-steps has a very significant consequence, the effect of which can be seen in experiments: the weights $\mu_{n}$ that are output with the samples are more accurate than if we for example relied on the stepsize that was used to advance the previous time-step.
In this method the implementation of the underlying Langevin integrator itself is as in the fixed-stepsize setting. There is no additional significant cost overhead above the usual cost of fixed stepsize integration if, for example, the BAOAB integrator is used for Langevin dynamics, since the force evaluation needed to update Z is already performed in the BAOAB iteration.
We note that initialization of Algorithm 1 requires a choice for $\zeta_{0}$. In our experiments we typically chose the initial conditions $(x,p)$ so that the force was small and thus used an intial value of $\zeta_{0}=0$. Since $\psi(\zeta_{0})=M$ for either of the given choices of the filter function, we have $\Delta t_{0}=M\Delta\tau$. It is expected that there will be some equilibration of the timestep control in the first few steps of integration.
3.2 Convergence and Order of Accuracy
The order of accuracy of a numerical method for SDEs can be studied in different contexts. In some disciplines, the quantity of interest is the strong accuracy, defined as the fidelity of the numerical solution to the solution of the SDE associated to some realization of the Wiener process. For sampling, the more relevant quantity is the accuracy of the approximation of the distribution generated by SDE solutions at a specified time (weak accuracy). An article of G. Vilmart [99, Proposition 6.1] provides a rigorous proof of the weak convergence of general splitting methods, which involve components which can be integrated exactly (in the weak sense) but which may contain multiplicative noise, as is the case for ZBAOABZ and other such integrators for SamAdams. Specifically we can state a theorem regarding finite time approximation based on this work (see Theorem 1).
Theorem 1.
Consider the system (10)-(13) and assume that $\psi$, $\sqrt{\psi}$, $\nabla U$, $g$ are $C^{6}$ functions with all partial derivatives bounded, further assume that $\psi>r$ for some $r>0$. Then consider splitting into A, B, O and Z components and generating a sequence of points $(x_{n},p_{n},\mu_{n})$ based on Algorithm 1 with any symmetric splitting with these components. Let $\chi(\cdot)\in C^{6}$ be an observable function of $x,p$ where all partial derivatives have polynomial growth, then for all $\Delta\tau k\leq T$
$\left|\frac{\mathbb{E}(\chi(x_{n},p_{n})\mu_{n})}{\mathbb{E}(\mu_{n})}-\frac{\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))}{\mathbb{E}(\psi(\zeta(n\Delta\tau))}\right|\leq C\Delta\tau^{2},$ (18)
where $C>0$ is independent of $\Delta\tau>0$ and $(X(\cdot),P(\cdot),\zeta(\cdot))$ is the solution to (10)-(12).
Proof.
We first remark that $\mathbb{E}(\mu_{n})>0$ and $\mathbb{E}(\psi(\zeta(n\Delta\tau)))>0$ for all $n\in\mathbb{N}$ due to the uniform bound assumption on $\psi$. Let $x=\mathbb{E}(\chi(x_{n},p_{n})\mu_{n})$, $y=\mathbb{E}(\mu_{n})$, $w=\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))$ and $z=\mathbb{E}(\psi(\zeta(n\Delta\tau))$. Then we have
$\left|\frac{x}{y}-\frac{w}{z}\right|$ $=\left|\frac{(x-w)z+w(z-y)}{yz}\right|\leq\frac{|x-w||z|+|w||z-y|}{|y||z|},$
and we have that $|x-w|\leq C\Delta\tau^{2}$ and $|z-y|\leq C\Delta\tau^{2}$ by [99, Proposition 6.1], then due to the uniform lower bound on $\psi$ we have the required result. ∎
Remark 1.
If initialized according to the invariant measure of (10)-(13) and if the process is ergodic we have that
$\frac{\mathbb{E}(\chi(X(n\Delta\tau),P(n\Delta\tau))\psi(\zeta(n\Delta\tau))}{\mathbb{E}(\psi(\zeta(n\Delta\tau))}=\mathbb{E}_{\pi_{\beta}}(\chi)$
for all $n\in\mathbb{N}$ following from Section 2.1.
The weak convergence can in principle be studied in the asymptotic sense as the time interval tends to infinity, i.e., we may consider the asymptotic evolution of the weak error. The challenge then is to establish the convergence rate in a suitable framework and to study the systematic bias that is introduced due to discretization. Theoretical study of the geometric convergence of Alg. 1 and the order of accuracy of the stationary distribution of the numerical method will be explored in future work, using techniques that are by now well developed in the setting of Langevin dynamics [11, 57, 34, 23, 68, 60].
It should be noted that ZBAOABZ does not have the additional desirable properties of BAOAB such as its property of quartic accuracy for configuration variables at high friction and its exactness for Gaussian targets [59], but the primary motivation for sampling methods is typically to provide stability and fast exploration of the state space. In settings where the energy landscape is complex, there can be considerably higher error due to lack of exploration than due to the bias arising from the numerical discretization. We illustrate this in the examples.
4 Adam-inspired Monitor Function and Sundman Transform
The time-rescaled Langevin dynamics (10)-(13) yields a family of samplers, each member specified by a particular choice of driving function $g(x,p)$ and Sundman transform kernel $\psi(\zeta)$. The choice of the monitor function, the auxiliary dynamics and the restriction function collectively decide the performance of the method, but have no impact on its theoretical foundation. In this section, we propose a choice that allows to adapt some of the advantages of the Adam optimizer to the realm of sampling.
In [22] it was demonstrated that the Adam optimizer can be interpreted as the Euler discretization of a certain system of ODEs, namely
$\mathop{}\!\text{d}x_{\tau}$ $=\frac{p_{\tau}}{\sqrt{\zeta_{\tau}+\epsilon}}\mathop{}\!\text{d}\tau,$ (19)
$\mathop{}\!\text{d}p_{\tau}$ $=-\nabla U(x_{\tau})\mathop{}\!\text{d}\tau-\gamma p_{\tau}\mathop{}\!\text{d}\tau,$ (20)
$\mathop{}\!\text{d}\zeta_{\tau}$ $=[\nabla U(x_{\tau})]^{2}\mathop{}\!\text{d}\tau-\alpha\zeta_{\tau}\mathop{}\!\text{d}\tau,$ (21)
with $x_{\tau},p_{\tau},\zeta_{\tau}\in\mathbb{R}^{d}$, $\alpha>0$, $\gamma>0$, and the algebraic operations are to be understood elementwise. To see how a discretization of this ODE leads to the Adam optimizer, see Appendix B and [50].
Assuming $\beta^{-1}$=0 in the time-rescaled process (10)-(13), one observes that the choices
$\psi(\zeta)$ $:=\frac{1}{\sqrt{\zeta+\epsilon}},$ (22)
$g(x,p)$ $:=\|\nabla U(x)\|^{2},$ (23)
mimic the Adam ODE (19)-(21).
The $\zeta$-dynamics (9) then becomes
$\mathop{}\!\text{d}\zeta_{\tau}=-\alpha\zeta_{\tau}\mathop{}\!\text{d}\tau+\|\nabla U(x_{\tau})\|^{2}\mathop{}\!\text{d}\tau.$ (24)
Similar to Adam, the solution $\zeta(\tau)$ then computes the weighted average of $\|\nabla U\|^{2}$ over recent history, as determined by the exponential weight $\alpha$, see (16)-(17) and the discussion in Appendix C.
The resulting dynamics differs in structure from that of Adam in two key ways:
  1. 1.
    In Adam, the choice is made to leave the ODE for the momentum untransformed. In other words, the Adam ODE (19)-(21) is an incomplete Sundman transform. Viewing the dynamics (19)-(21) in real time $t$ leads to
    $\mathop{}\!\text{d}x_{t}$ $=p_{t}\mathop{}\!\text{d}t,$
    $\mathop{}\!\text{d}p_{t}$ $=-\frac{1}{\psi}\nabla U(x_{t})\mathop{}\!\text{d}t-\frac{\gamma}{\psi}p_{t}\mathop{}\!\text{d}t,$
    which implies that Adam uses an effective configuration-dependent momentum evolution. By contrast, our method adapts the timescales of position and momentum components in a symmetric way.
  2. 2.
    Adam uses a vectorial $\zeta\in\mathbb{R}^{d}$, which in the language of time-rescaling corresponds to an individual adaptive stepsize per degree of freedom. The squared Euclidean norm in (23) is replaced by an elementwise squaring in (21). Incorporating such individual timesteps in a sampling context introduces additional complication in both the theoretical foundation and practical implementation, and is left for future work.
While we could employ the Sundman kernel (22), we follow the idea of [43, 61] and introduce a filter in $\psi$ in order to restrict the value to a specified interval. Two (similar) choices for Sundman transformation are
${\psi}^{(1)}(\zeta)=m\frac{\zeta^{r}+M}{\zeta^{r}+m},\hskip 21.68121pt{\psi}^{(2)}(\zeta)=m\frac{\zeta^{r}+M/m}{\zeta^{r}+1},$ (25)
for two constants $0<m<M<\infty$. We see that, for either choice, ${\psi}(0)=M$ and $\psi(\infty)=m$ such that $m$ and $M$ serve as bounds on the Sundman transform kernel and hence on the adaptive stepsize, which then satisfies $\Delta t\in[\Delta t_{\rm min},\Delta t_{\rm max}]$, where $\Delta t_{\rm min}=m\Delta\tau$ and $\Delta t_{\rm max}=M\Delta\tau$. In particular, $M/m$ gives the maximum factor by which the timestep can be dilated relative to the minimum stepsize. Using $\psi^{(1)}$ or $\psi^{(2)}$ can improve stability compared to (22) and allows the user to exert more control on the effective stepsizes $\Delta t$. The power $r>0$ can additionally be used to adjust the dependency to influence the distribution of stepsizes used in simulation. We note that for $\zeta\gg 0$, [^2] To see this, divide numerator and denominator of $\psi^{(2)}$ by $\zeta^{r}$ and expand in a geometric series.
$\psi^{(2)}(\zeta)\sim m+(Mm-m^{2})/\zeta^{r}.$
Hence for $m\approx 0$, $Mm=1$, we see that this is asympotically related to $\zeta^{-r}$ for large $\zeta$, i.e., where the smaller steps are needed. For further discussion of how the $\zeta$-dynamics (24) together with the transform kernel $\psi$ influences the stepsize adaption, we refer to Appendix C.
An Adam-like choice of $g$ is $g(x,p)=\|\nabla U(x)\|^{2}$, i.e., monitoring the force norm. The closest fidelity to (22) is thus found with $r=1/4$ in the Sundman transformation, since scaling the full vector field by $a$ can be related to scaling half the system only by $a^{2}$. For flexibility we propose to use $g(x,p)=\Omega^{-1}\|\nabla U(x)\|^{s}$, i.e. monitoring the force norm raised to some possibly fractional, positive power $s$ and scaled by a normalization factor $\Omega^{-1}$. Effectively, this choice combined with raising $\zeta$ to the $r$th power in the restriction function is tantamount to controlling the timestep based on the $r\cdot s$ power of the gradient norm.
Where it is not a physical parameter of the modelling task, the temperature $\beta^{-1}$ allows the method to behave more like an optimizer ($\beta^{-1}\approx 0$) or a sampling scheme ($\beta^{-1}\rightarrow 0$). In a Bayesian sampling context, $U(x)$ is the negative log-posterior and we may take $\beta^{-1}=1$ to sample the posterior or $\beta^{-1}<1$ to implement annealed importance sampling [72]. Due to the various points mentioned above, SamAdams with $s\cdot r=1/2$ does not strictly reduce to Adam for $\beta^{-1}\rightarrow 0$. In our experiments we found that the optimal choices of $r$ and $s$, as well as other coefficients, were problem-class dependent, but within a specific class of models those selections were relatively easy to decide.
4.1 Other Monitor Functions
The choice to make SamAdams relate to the Adam optimizer is not unique: possible alternatives include basing $g$ on only the prior in a Bayesian sampling setting or basing $g$ on some subset of the variables (which are known to exhibit high levels of variability). It would also be possible to introduce higher derivative information derived from the potential energy function such as the trace or determinant of the Hessian matrix. We have not substantially explored such options. For noisy gradient evaluations, one might also design a monitor function based on the estimation of gradient noise such that stepsizes are decreased whenever the injected gradient noise is large. This idea is motivated in Appendix D, which shows results for logistic regression with stochastic gradients, during which the adaptive stepsize strongly reacts to the employed batch size.
5 Numerical Experiments
5.1 Asymmetric Double Well
We first demonstrate the stepsize adaptation of SamAdams on a one-dimensional double well problem where one well is much narrower than the other. The potential is given by $U(x)=\frac{b}{L}(x+1)^{2}(x-L)^{6}$ with $b=1.5$ and $L=2$. We pick a small temperature $T\equiv\beta^{-1}=0.4$ to make the transition across the barrier reasonably rare. In the narrow well, a Langevin path will encounter greater forces, causing SamAdams to decrease its stepsize $\Delta t$ accordingly. Fig. 3 (top) shows the potential and the density on the left, and the $x$- and $\Delta t$-values of a single trajectory of SamAdams on the right. We clearly see that occupation in the narrow well (negative $x$-values) leads to a restriction of the stepsize $\Delta t$ to small values.
Figure 3: Sampling experiments on a 1D toy model (see text). (a) Potential and density for employed temperature $T=0.4$. (b) $x$-coordinate and adaptive stepsize $\Delta t$ for SamAdams along a single trajectory. The black dashed line gives the value of virtual stepsize $\Delta\tau$ which is adaptively increased or reduced to yield the real stepsize $\Delta t$. (c) Mean absolute errors of two observables, the $x$-coordinate and the occupation frequency of area $x<0.5$ against (mean) stepsize $\Delta t$. The different values for SamAdams were obtained by varying $\Delta\tau$ from 0.03 to 0.2. (d) $\Delta t$ histograms for SamAdams run at three different $\Delta\tau$. Other hyperparameters: $\gamma=\alpha_{1}=\alpha_{2}=1$, $m=0.1$, $M=10$, $r=0.25$, $s=2$.
The lower left panel of Fig. 3 shows absolute mean errors of two observables, the $x$-coordinate and indicator function of the domain $x<0.5$ (which roughly corresponds to the occupation probabilities of the barrier and the narrow well, respectively), against stepsize (mean stepsize for the adaptive scheme). We first ran SamAdams for different values $\Delta\tau$, measured the mean adaptive stepsizes $\langle\Delta t\rangle$ for each of these runs, and then ran BAOAB at stepsizes fixed to these values to obtain canonical averages at the same mean stepsize and compute cost. Each point in the figure was generated by averaging over 300 independent trajectories for $5\cdot 10^{7}$ iterations (discarding the first 100,000 iterations as burn-in). The ground truths $\langle x\rangle$ and $P(x<0.5)$ were obtained via numerical quadrature. We also plot vertical lines denoting the maximum (mean) stepsize at which the corresponding algorithm became unstable in at least one of the 300 trajectories (i.e., the stability threshold). We see that BAOAB does significantly worse than SamAdams, only reaching similar accuracies for the smallest stepsize examined. SamAdams’ performance barely depends on $\langle\Delta t\rangle$ until very close to its stability threshold. We also observe that it is able to use larger steps than BAOAB.
The bottom right of Fig. 3 shows the $\Delta t$ histograms for three of the SamAdams runs. They have a bimodal structure, as might be expected from an asymmetric double well. All three of them are able to use stepsizes larger than the stability threshold of BAOAB, supporting the idea that the stability threshold for fixed-stepsize schemes depends on local variation of the loss landscape, rather than on the landscape as a whole. As for the star potential of Fig. 1, it is enough to use small stepsizes in critical areas. Note how one of the $\Delta t$ histograms has a mean of $\langle\Delta t\rangle=0.216$, which is 31% larger than BAOAB’s threshold, which together with the low observable errors even at that stepsize implies a substantial increase in computational efficiency compared to BAOAB. The ability to use larger stepsizes (in some cases much larger stepsizes) than constant stepsize schemes while preserving sampling quality will also be observed in the examples of the next subsection.
5.2 Planar Systems
In the introduction, we already mentioned an example involving the “star potential”, which has narrow corridors that can be difficult to sample efficiently. Here we consider several other examples of 2-degree of freedom problems and explore the accuracy and stability of the new method in comparison with fixed stepsize integration.
5.2.1 Neal’s Funnel
We consider a 2D variant of Neal’s funnel [72] (a 9d version will be taken up in the following subsection). The potential energy function is
$U_{\rm Neal}(x,\theta)=\frac{x^{2}}{2e^{\theta}}+\frac{\epsilon}{2}(x^{2}+\theta^{2}).$
Our goal is to sample the canonical distribution at temperature $T=1$. We used $\gamma=5$ and discarded $10^{5}$ steps as burn-in (equilibration). The initial position was taken to be $(0,5)$ in a relatively flat zone, and we set initial momenta and $\zeta_{0}=0$.
Because canonical sampling may be intractable due to unconstrained domains, one often incorporates a term in the form of a prior (and associated potential) to maintain confinement of solutions; we have done this here by using a simple harmonic restraint.
In the funnel problem the domain shrinks to a narrow neck as $\theta\rightarrow-\infty$. This creates a numerical challenge as the trajectory rattles back and forth against the walls of the channel; for a fixed moderate stepsize, at some point the solution will become unstable and jump out of funnel. In practice these escaping trajectories often re-enter the larger domain (with $\theta>0$) and re-equilibrate at the target temperature, but the unstable behavior can damage the computation of observables. An example of this type of behavior can be seen in Fig. 4 where a trajectory is shown together with the evolution of kinetic temperature and average potential. As we can see, the two observables are severely degraded when the instability is encountered which would ultimately be seen as poor convergence. In the right panel of Fig. 4, we also see what happens when the stepsize is halved ($\Delta t=0.02$). The instabilities are still very much in evidence if the trajectory is long enough (here $N=10^{8}$). The stepsize would need to be below $\Delta t=0.01$ to completely eliminate the instability. [^3] Discrete Langevin trajectories for potentials that are not globally Lipschitz are inherently unstable, due to the use of normally distributed random variables (see [95, 67] for some discussion); that said, in our experience, for typical systems the frequency of long excursions in fixed stepsize trajectories decreases rapidly as the stepsize drops below a certain well defined stability threshold.
Figure 4: (left and center) A BAOAB trajectory with stepsize $\Delta t=0.04$ shows an unstable evolution in $10^{6}$ steps. The descent into the funnel leads to a spike in both kinetic temperature and mean potential energy. Although shortlived, this type of event can, as here, corrupt long term averages. (right) At longer times, these instabilities are inevitable, for stepsizes above or equal to $\Delta t=0.015$.
By contrast, SamAdams (ZBAOABZ) produces reliable, stable trajectories with much larger mean stepsize than any fixed stepsize method. In Fig. 5 we show a trajectory with mean stepsize $\langle\Delta t\rangle=0.16$ ($N=10^{7}$). (The details of stepsize variation are as follows: $\Delta t_{\rm min}=10^{-4}$, $\Delta t_{\rm max}=0.6$, $\Delta t_{0}=0.6$, $\alpha=0.1$, $r=0.5$ and $g(x,p)=\|\nabla U\|$. We used the second form of the filter function $\psi=\psi^{(2)}$ in all the 2d examples.)
A histogram computed using the samples obtained from SamAdams is virtually identical to the target distribution (Fig. 6), despite requiring around 10% of the computational effort needed if the corresponding fixed-stepsize method was used. Observables, e.g. the kinetic temperature or the mean potential energy as shown in Fig.5 are approximated to three significant digits.
Figure 5: A SamAdams trajectory with mean stepsize $\langle\Delta t\rangle=0.16$ corrects the instability of the fixed stepsize method. The kinetic temperature and potential energy average converge to three significant digits of accuracy.
Figure 6: Here we compare the (weighted) histogram of solution data to the actual canonical distribution for the SamAdams trajectory with $\langle\Delta t\rangle=0.16$. The distributions are visually very similar.
5.2.2 Entropic Barrier
Rare event sampling problems in many fields can be characterized by low energy basins connected by thin channels. Diffusion through the narrow corridors typically requires small stepsizes since too-large stepsizes either lead to expulsion from the corridor or numerical instability. An entropic barrier problem was constructed to illustrate the challenge of such tasks. The potential function is
$U_{\rm channel}(x,y)=\frac{y^{2}}{1+10x^{4}}+0.001(x^{2}-9)^{2}.$
We simulate this with temperature set to $0.05$ to create a challenging test. We used friction $\gamma=5$ in this example which gave approximately optimal results.The initial point was taken at $(x,y)=(3,0)$ near the minimum on the right side, and the initial momenta were set to zero. The SamAdams parameters were like in the previous example ($r=0.5$, $\alpha=0.1$), but we set $\Delta t_{\rm min}=0.0001$, $\Delta t_{\rm max}=0.5$ $\Delta t_{0}=0.5$.
Figure 7: Comparison of trajectories obtained using fixed and variable stepsize. (a) BAOAB with fixed stepsize $\Delta t=0.1$ converges as expected, with around 2-3 barrier crossings per 1M steps. (b) BAOAB becomes unstable above $\Delta t=0.15$ and at $\Delta t=0.2$ shows no diffusion over the barrier. (c) A variable stepsize trajectory with $\langle\Delta t\rangle=0.356$, restoring the performance of small fixed stepsize.
Figure 8: Large variable stepsize results ($\langle\Delta t\rangle=0.356$). In the left figure, a histogram of the trajectory data (weighted by the time-rescaling used to generate the data) is shown. This can be compared to the exact distribution (from the probability density function) shown in the central panel. Finally the actual stepsize distribution is given in the right panel and shows that a vanishingly small number of steps require a small (below 0.2) stepsize. Not visible in the histogram is the fact that a very small number of steps used a stepsize below 0.01.
We found that the fixed stepsize integrator was stable up to a maximum stepsize of $\Delta t\approx 0.2$, for many trajectories, but above $\Delta t=0.15$ there is a lot of error in the narrowest part of the channel and the number of full crossings from one basin to the other is significantly reduced. Fixed stepsize trajectories with $\Delta t=0.1$ and $\Delta t=0.2$ are shown in Fig. 7. Also shown is a typical SamAdams trajectory with mean stepsize $\langle\Delta t\rangle=0.356$.
In the left panel of Fig. 8 we show a (weighted) histogram of the computed states for $\langle\Delta t\rangle=0.356$ which can be compared with the exact distribution in the central panel. In this example, with $10^{8}$ steps, the kinetic and configurational temperatures (observables which should each average to the target temperature [59]) are accurate to within 1% ($T_{\rm kin}=0.04947$, $T_{\rm conf}=0.04954$). Of particular interest is the fact that the stepsize distribution shown at right in Fig. 8 has barely any mass below $\Delta t=0.2$, meaning that the small stepsizes are only needed at very rare instances of barrier crossing (precisely in a small neighborhood of the origin).
5.2.3 Beale Potential
In this 2D model there are again two basins, but they have a complicated curved shape. The rarity of transitions is actually still more extreme than in the entropic barrier problem. The two wells have very different depths and shapes. Transitions between the wells happen in a narrow region around the origin, although we have noticed that numerical error can either eliminate corridors or create new ones. With a 6th order exponential confinement term, the Beale potential takes the form
$U_{\rm Beale}(x,y)$ $=$ $(1.5-x+xy)^{2}+(2.25-x+xy^{2})^{2}+(2.625-x+xy^{3})^{2}$
$\hskip 21.68121pt+0.3\exp\left(0.00001(x^{6}+y^{6})\right).$
We used a temperature of $T=3$ to obtain sufficient barrier transitions. The initial point was $(x,y)=(3,0)$, with zero momenta.
For fixed stepsize integration, the step needs to be below about 0.003. Running at or above this threshold almost all runs of $10^{8}$ steps result in failure. At the small value $\Delta t=0.0025$, $10^{6}$ samples are not enough to cover the distribution, as shown in Fig. 9. By contrast (see the right panel of Fig. 9) we were able to use SamAdams (ZBAOABZ) reliably with a mean step of $\Delta t=0.022$ which indicates a stability improvement of nearly an order of magnitude. While there is some visible error in the histogram for SamAdams (Fig. 9) it is much less than for BAOAB operating at its much smaller stable stepsize.
Figure 9: For the Beale potential, we show the probability distribution at the left and the histogram (center) for a fixed stepsize BAOAB run with $\Delta t=0.0025$ (larger stepsizes are unreliable), $N=10^{6}$ steps. Finally, at right we see a histogram obtained from a SamAdams trajectory $\langle\Delta t\rangle=0.022$, $N=10^{6}$ steps. Parameters: $\Delta t_{\rm min}=0.001$, $\Delta t_{\rm max}=0.1$, $\Delta t_{0}=0.1$ $r=0.5$, $\alpha=1$, and monitor function $g(x,p)=\|\nabla U(x)\|$.
We also perform more extensive runs to resolve the bias in canonical averages. The procedure is similar to the case of the 1D toymodel in Sec. 5.1. To obtain different $\langle\Delta t\rangle$ for SamAdams, we varied $\Delta\tau$, keeping everything else fixed, and then ran BAOAB at fixed stepsizes set to the obtained $\langle\Delta t\rangle$ values for comparison. We averaged over 200 independent trajectories, where the number of iterations scaled with $\Delta\tau$ (we used $10^{8}$ iterations at $\Delta\tau=0.001$, and varied $\Delta\tau$ from 0.00025 to 0.06). The computational cost of each BAOAB simulation was the same as for the corresponding SamAdams run. Fig. 10 shows the bias against (mean) stepsize $\Delta t$ for the two temperatures and coordinates (the ground truths for the latter were obtained via numerical quadrature). While SamAdams and BAOAB show comparable errors for small stepsizes, BAOAB becomes unstable at $\Delta t=0.0025$. SamAdams can still be run at $\langle\Delta t\rangle=0.01$ without any decline in accuracy. If a modest decline in accuracy is acceptable, it can be run at substantially larger steps. In fact, while the BAOAB curves in Fig. 10 stop at the last stable stepsize, SamAdams remained stable until the last examined $\Delta\tau$.
Figure 10: Absolute errors of the means of four observables on the Beale potential against (mean) stepsize $\Delta t$. The ground truths for the coordinates were obtained via numerical quadrature. Hyperparameters: $\gamma=1$, $T=3$, $m=0.1$, $M=10$, $r=0.25$, $\alpha=1$, $g(x,p)=0.1\|\nabla U\|^{2}$. See text for more information.
5.3 Higher-dimensional Models
We now explore several higher-dimensional cases using fixed stepsize Langevin as well as the SamAdams scheme. Our goal here is to show, using model systems, that there is a good basis for believing the method will be effective for applications in statistics and machine learning. SamAdams has several parameters ($\alpha$, $m$, $M$, $r$, and indeed the monitor function $g$) that can be tuned to adjust the performance of the method. In practice we have found that some attention needs to be paid to their selection to obtain optimal results.
5.3.1 Neal’s funnel (multi-dimensional case).
The 2D funnel example of the last subsection is a simplified version of a model with multiple latent variables that was originally proposed as a surrogate for problems in Bayesian hierarchical modeling with nonlinear dependencies [72]; for a related model arising in ecology, see [69]. We test SamAdams on the original 9-dimensional model modified with a confining prior with a large variance ($\sigma_{x}^{2}=20$) to ensure sufficiently frequent trips into the funnel neck. We used the following setup in our simulations:
$\theta\sim{\mathcal{N}}(0,3),\hskip 14.45377pt\left.x_{i}\right|\theta\sim{\mathcal{N}}(0,\exp(\theta))\times\mathcal{N}(0,\sigma_{x}^{2}),$
with p.d.f.
$p(\theta,x_{1},x_{2},\ldots,x_{n})=\frac{1}{\sqrt{(2\pi)^{n}\cdot 6}}\exp\left(-\frac{\theta^{2}}{6}\right)\cdot\exp\left(-\frac{n}{2}\theta-\sum_{i=1}^{n}x_{i}^{2}\left(\frac{1}{2e^{\theta}}+\frac{1}{2\sigma_{x}^{2}}\right)\right).$
Figure 11: Comparison of trajectory graphs in the $\theta,x_{1}$ projection obtained for different methods and stepsizes. Top row: fixed stepsize BAOAB runs with $\Delta t$ increasing left to right until explosion is encountered at $\Delta t=0.1$. The lower row shows the results of different variable stepsize runs using SamAdams. All of these solutions are usable for sampling purposes, although the last one shows some thickening in the $x$ direction.
We examined the stability of fixed stepsize (BAOAB) Langevin dynamics by running trajectories with a series of increasing timesteps $\Delta t=0.04,0.06,0.08,0.1$. The system was initialised with a large enough value of $\theta$ ($\theta_{0}=5$) and all $x_{i}=0$ so as to avoid any initial high gradient. The graphs in the $\theta,x_{1}$ projection are shown in the upper row of Fig. 11. All simulations involved $N=10^{7}$ steps. We then ran SamAdams with various $\Delta\tau$, generating the figures shown in the second row, also using $10^{7}$ steps for each. Parameters of SamAdams simulations: $m=0.01$, $M=1.0$, $\alpha=1$, and $r=1$, with monitor function
$g(x,p)\equiv\frac{1}{100}\|\nabla U(x)\|$
and filter function $\psi=\psi^{(1)}$. Here $m,M$ define a range of 100 in stepsize, smallest to largest. We label these runs by the mean stepsize used, which is reciprocally related to the computational work to reach a given fixed time. As we can see, all the SamAdams simulations are stable and accurate representations of the true SDE sampling trajectory. Given the mild confining potential, the excursions to large positive $x$ in the fixed stepsize trajectories at stepsize greater than $0.04$ represent high energy (far from equilibrium) events; in a more complicated model those trajectories would lead to blow-up, or, perhaps more seriously, subtle degradation of cumulative averages; here the confining potential quickly returns them to the domain of interest. By contrast, SamAdams is stable and retains the trajectories within a similar size and shape region of the $\theta,x_{1}$ plane, except for a slight increase in the width of the sampled region at the largest mean stepsize ($\langle\Delta t\rangle=0.1762$).
One way to read the results of Fig. 11 is that the largest stable mean stepsize of SamAdams is approximately four times the largest stable stepsize for Langevin dynamics with fixed stepsize.
In Fig. 12, we show a SamAdams trajectory with the points colored by the stepsize used. We see that the smallest stepsizes are only used at the narrowest part of the funnel neck and the stepsize quickly resets as we leave these locations. Stepsize distributions are shown in the right panel of Fig. 12, with the mean stepsizes used in the four SamAdams runs of Fig. 11. The parameters $m$ and $M$ partly govern the shape of these distributions, which are also defined by features of the problem itself.
Figure 12: Left: points along a trajectory of the 9D Neal funnel (here for $\langle\Delta t\rangle=0.1367$) are colored by stepsize used. Right: the actual stepsize distributions for four SamAdams trajectories are shown; note that the stepsize scale at right is logarithmic, meaning that there is less density associated to the smaller stepsizes than is apparent from the areas of the respective bins.
An important question is whether the time-series corresponding to different timesteps explore the space with similar efficiency. We settle this for the 9D funnel example in Fig. 14 where we see that relative to the elapsed time, the trajectories diffuse at similar rates. [^4] Computing autocorrelation functions and ESS requires first interpolating the non-uniformly spaced time-series data to a uniform mesh; this interpolation can be avoided when computing expectations by using the method of Section 2.1. Effective sample sizes per sample [18] are as shown in Table 13. These results indicate that the trajectories are similar in terms of the rates of exploration.
Mean log posterior values (from 10M step runs) for the different stepsizes are also given in the table and suggest that bias may be becoming more noticeable at the larger stepsize. The result $-10.46$ for mean log posterior is accurate, verified using a small stepsize of $0.01$ and 1 billion steps. Note that with small stepsize $0.01$, fixed stepsize runs of 10M steps as in the table generate similar size errors (in this case due to Monte Carlo error due to the higher correlation of samples) to the largest variable stepsize integration (in that case, due to sampling bias), thus demonstrating a clear trade-off between sampling error and bias. Accuracy (and ESS) appear to fall off at the largest fixed stepsize with $\Delta t=0.02$ being approximately optimal. This can be compared to SamAdams with a mean stepsize of $\langle\Delta t\rangle=0.0998$, indicating an improvement of over 400% in sampling efficiency without any reduction in accuracy.
fixed $\Delta t$ variable $\Delta t$
mean $\Delta t$ 0.01 0.02 0.04 0.0664 0.0998 0.1336 0.1676
ESS/sample (×$\times 10^{-4}$) 6.6 6.2 5.6 6.2 6.3 6.3 6.5
mean log posterior($10^{7}$ samples) -10.55 -10.47 -10.42 -10.46 -10.46 -10.50 -10.53
Figure 13: Table showing the effective sample size per sample and the expected log posterior, for different variable stepsizes.
Figure 14: Autocorrelation functions with respect to iteration number generated by different mean stepsizes. Left: autocorrelation of $\theta$; right: autocorrelation of $x_{1}$. For each variable, the decay rates corresponding to different schemes are similar, regardless of the stepsize (and are similar for both the fixed and variable stepsize runs).
5.3.2 Auxiliary Variable Control Dynamics
Until now we have not yet explored the role of the hyperparameter $\alpha$ in (24) nor the influence of the scale of the monitor function $g(x,p)$. In the case of $g(x,p)=\Omega^{-1}\|\nabla U(x)\|^{r}$ (our choice for all of our experiments in this article), the scaling factor $\Omega$ governs how strongly $\zeta$ reacts to a given force value $\nabla U(x)$. Since the range of typical values of $\nabla U$ strongly depends on the problem at hand, the scaling parameter $\Omega$ is useful to adjust the range of $g(x,p)$ and hence the range of $\zeta$ and $\Delta t$. For the one- and two-dimensional examples considered thus far, setting $\Omega=1$ serves as a good starting point, whereas on the classification tasks of the the next section we use $\Omega=N_{\text{D}}$ with $N_{\text{D}}$ the size of the dataset. Since the force $\nabla U$ is written as a sum over the model likelihoods of all data points, this choice for $\Omega$ encourages $\zeta$-values close to the fixed point of the Sundman transform $\psi$. The hyperparameter $\alpha$ governs the damping of the $\zeta$-dynamics as well as its moving average behavior, i.e., how quickly past force values are ’forgotten’ (see Appendix C for more details).
As we have shown thus far, one fundamental benefit of the adaptive stepsize is the ability to use overall large stepsizes while using small stepsizes selectively. If the resulting $\langle\Delta t\rangle$ is larger than the maximum stable stepsize of a conventional scheme (assuming the accuracies remain acceptable), SamAdams leads to an improvement of stability and computational efficiency. An important question is then how difficult it is to tune the SamAdams-specific hyperparameters $\alpha$ and $\Omega$. To examine this, we chose the example of the star potential. We create a grid of different $(\alpha,\Omega)$-values and run SamAdams trajectories on for each grid point. For all ($\alpha,\Omega$), we experimentally find the stability threshold of SamAdams, i.e., the largest $\langle\Delta t\rangle$ that still leads to a stable trajectory. We do this by running 100 independent trajectories for successively increasing values of $\Delta\tau$, the base stepsize, until the trajectory becomes unstable. We vary $\Delta\tau$ in $[0.04,0.08]$, a range in which the resulting $\langle\Delta t\rangle$-values become roughly comparable to the stability threshold of BAOAB for most ($\alpha,\Omega$)-combinations  [^5] BAOAB’s threshold was found to be $\Delta t=0.01275$, defined as the smallest stepsize for which 100 independent trajectories remain stable for roughly $5\cdot 10^{6}$ iterations.. The number of iterations per tested $\Delta\tau$ is scaled via $N=N_{0}\Delta\tau_{0}/\Delta\tau$ with $N_{0}=2\cdot 10^{6}$ and $\Delta\tau_{0}=0.06$. If an $(\alpha,\Omega)$-combination leads to unstable trajectories even at the smallest $\Delta\tau$ tested, we deem the combination ’highly unstable’. The results are shown in Fig. 15.
Figure 15: SamAdams stability thresholds $\langle\Delta t\rangle_{\text{max}}$ in dependency of $\alpha$ (attack rate) and $\Omega$ (scale coefficient of force norms) in the case of the star potential. Left: Plain $\langle\Delta t\rangle_{\text{max}}$. A well defined optimal zone appears around $\alpha=1$, $\Omega=100$. Right: Fraction of SamAdams threshold and BAOAB’s threshold $\Delta t^{\text{BAOAB}}_{\text{max}}$. The brown areas show parameter regions where BAOAB is more stable than SamAdams, the purple area regions where SamAdams is more stable. The red areas denote ’highly unstable’ regions (see main text). Other hyperparameters: $T=\gamma=1$, $m=0.1$, $M=10$, $r=0.25$, $s=2$, transform kernel $\psi^{(1)}$.
From the left-hand figure, we see how SamAdams’ stability threshold varies with $(\alpha,\Omega)$, with an optimum region visible for $\alpha\simeq 1$ and $\Omega^{-1}\simeq 0.01$. The highly-unstable area (red) is mainly due to too large values of $\alpha$. As shown in Appendix C, too large $\alpha$ will lead to $\Delta t=M\Delta\tau$. For our choice $M=10$ and the smallest $\Delta\tau$-value tested, $\Delta\tau=0.04$, this would correspond to a constant-stepsize scheme running at stepsize 0.4, more than 30 times larger than BAOAB’s stability threshold. Note that even for these highly unstable ($\alpha,\Omega)$-values, one could still enforce stability by picking $\Delta\tau$ and $M$ such that $M\Delta\tau$ is smaller than BAOAB’s threshold. On the right-hand side in Fig. 15, we plot the same grid but color according to the fraction of SamAdams stability threshold and BAOAB’s threshold. We observe that there is a large area in parameter space (shaded purple) in which SamAdams is more stable than BAOAB, with a remarkable improvement of up to 300%.
While these experiments show the need to pick admissible $(\alpha,\Omega)$-values, they also demonstrate that there is a wide range of settings in which SamAdams is more stable than BAOAB, reducing the need for hyperparameter fine-tuning.
5.3.3 MNIST Image Classification on an MLP
We next apply our new scheme to the MNIST digit classification dataset [28], a standard benchmark of computer vision applications. It consists of 60,000 training and 10,000 test examples of handwritten digits, stored as $28\times 28$ pixel grayscale images. As is customary, we normalize the data to mean 0.5 and standard deviation 0.5. We use a multi-layer perceptron (MLP) with two hidden layers with 800 and 300 nodes, respectively. While conventional neural network training is typically done with small batches, we use large batch sizes in our sampling experiments for now to prevent additional gradient noise from obscuring the effect of the (already noisy) stepsize adaptation. (For a preliminary study of the effect of varying batch size, refer to Appendix D.) Here we fix $B=10,000$. It is not our aim to demonstrate state-of-the-art performance for image classification of our new scheme. Rather, we want to illustrate the potential benefits of using adaptive-stepsize methods when exploring neural network loss landscapes compared to constant-stepsize methods. For the experiments in this and the next subsection, we use tranform kernel $\psi^{(1)}$ with $m=0.1$, $M=10$, $r=0.25$, and sample at $T=\gamma=1$. All trajectories are initialized through PyTorch’s default (the momenta being drawn from their invariant Gaussian measure). The monitor function is taken as $g(x,p)=N_{D}^{-1}\|\nabla U(x)\|^{2}$ with $N_{D}=60,000$ the number of examples in the training dataset. $\zeta$ is initialized to $g(0,0)$. Fig. 16 shows typical trajectories for BAOAB and SamAdams, where the former was run at the mean adaptive stepsize used by the latter. We observe a large loss spike in the constant-stepsize scheme which is not present for the adaptive scheme. Before and after the spike the dynamics seem to align to high degree, which implies that it is indeed the force-sensitive stepsize adaptation of SamAdams that prevents the spike from forming.
Figure 16: Single trajectory results for an MLP on MNIST. Top panel: training loss; middle panel: $\zeta$ dynamics; lower panel: adaptive stepsize $\Delta t$. BAOAB was run at the mean adaptive stepsize $\langle\Delta t\rangle$ of SamAdams. In this example we chose $\Delta\tau=0.0002$, and $\alpha=50$.
This is also clearly evident from the sudden increase of $\zeta$ and corresponding decrease of $\Delta t$ at that point. Observe also that we pick a comparatively large value of $\alpha$ on this example, which before and after the spike leads to a continuous increase of the stepsize because the forces in these regions are sufficiently small.
While even the constant-stepsize scheme is able to recover from the instability, the appearance of spikes like this can lead to decreased performance in actual posterior sampling experiments, in which one averages over many different trajectories. To demonstrate this point, we draw 100 independent trajectories for both SamAdams and BAOAB, and average the resulting final accuracies. For SamAdams we use the same settings as in Fig. 16, i.e., $\Delta\tau=0.0002$ and $\alpha=50$. For BAOAB, we perform two experiments. In the first round, we set its learning rate to the mean of the stepsizes adopted by SamAdams, pooled from all 100 trajectories. In the second round, we take the mean of the pooled stepsizes again, but only consider the first 10% of $\Delta t$ samples adopted by SamAdams on each trajectory. The latter setup represents the idea that, for a fair comparison, the choice of stepsize for BAOAB should only be allowed to use information of the early stage of the SamAdams run. In contrast, using SamAdams’ mean adaptive stepsize uses information of all the stepsizes used by SamAdams, i.e., knowledge of the force evolution across the whole trajectory. This information would usually not be available to someone wanting to set the stepsize of a constant-stepsize scheme, so using the mean adaptive stepsize for the BAOAB runs gives the benefit of the doubt to BAOAB. At the same time, using the mean adaptive stepsize for BAOAB is the choice that leads to similar computational cost when run for the same number of iterations, which can also be the seen as the basis for a ’fair’ comparison. The final mean train and test accuracies together with 95%-confidence intervals are given by Table 1.666Note: these runs were performed without time averaging. We observe that SamAdams significantly outperforms both BAOAB setups.
ZBAOABZ BAOAB (mean of all $\Delta t$) BAOAB (mean of first 10% of $\Delta t$)
Train Accuracy (%) 94.0$\pm$0.1 92.3 $\pm$ 0.7 92.9 $\pm$ 0.5
Test Accuracy (%) 93.6$\pm$0.1 91.0 $\pm$ 0.7 91.5 $\pm$ 0.6
Table 1: Final accuracies averaged over 100 independent trajectories. Mean values and 95% confidence intervals. For the hyperparameters used and a description of the difference of the two BAOAB runs, see main text.
5.3.4 MNIST Image Classification on a CNN
Since image classifiers usually adopt convolutional neural networks (CNN) rather than fully connected ones, we repeat the experiment from the previous section on a simple CNN using three convolutional layers. The network architecture is given in Appendix E. Unless explicitly restated, the hyperparameters are the same as in the previous section. Fig. 17 (left) shows the results of a single SamAdams trajectory compared to three different BAOAB trajectories, each one corresponding to a different stepsize (the smallest stepsize adopted by SamAdams, the mean stepsize, and the largest stepsize). SamAdams outperforms all BAOAB runs in terms of loss, train, and test accuracy. In particular, the two BAOAB runs at larger stepsizes become unstable during the early epochs and fail to train completely. The smallest BAOAB stepsize leads to reasonable results, but the convergence speed as measured in number of epochs (i.e., compute time) is substantially smaller than for SamAdams, implying enormous computational speed-ups when using the latter. Fig. 17 (right) shows the evolution of $\zeta$ and $\Delta t$ of the SamAdams run compared to the loss. One observes how the algorithm reacts to the instabilities visible in the loss during the early training phase by rapidly reducing the stepsize $\Delta t$, well below the level of the learning rates used by the two unstable BAOAB runs. Only the BAOAB run using the smallest $\Delta t$ adopted by SamAdams remained stable, implying that small stepsizes are necessary to make it through the early stage of training. However, from epoch 5 onwards SamAdams slowly increases its stepsize again until it stabilizes at $\sim 0.0011$, more than three times the size of the stepsize used by the stable BAOAB run. This explains SamAdams faster convergence in loss and accuracy. From looking at the loss curves and the obtained $\Delta t$ by SamAdams, it seems like the trajectories start on a plateau (allowing for a rapid increase in $\Delta t$ at the very beginning), then descend down through an irregular landscape (leading to rapid damping in $\Delta t$ and breakdown of two of the three BAOAB runs), then reach a widened basin (allowing for moderate increase of $\Delta t$ again). This challenges the conventional wisdom in deep learning to use large learning rates at the start of training and successive decrease of learning rates during later phases [24, 4, 26]. Potentially, one could use adaptive stepsize schemes in the place of conventional learning rate schedulers, an idea the exploration of which we leave for future work.
Figure 17: Training of a CNN on MNIST. Left: Train loss, train and test accuracies. BAOAB was run at three different stepsizes: the smallest, largest, and mean stepsize used by SamAdams. Right: SamAdams results for loss (same as on the left), $\zeta$, and $\Delta t$. The dashed lines correspond to the stepsizes used by BAOAB. Hyperparameters: $\Delta\tau=0.002$, $\alpha=500$.
We now examine accuracies obtained by the two schemes via posterior sampling and averaging. Similar to the previous section, we run 100 independent trajectories, initialized as before. The hyperparameters are the same as in Fig. 17. Each trajectory is run for 60 epochs where the train and test accuracies are computed every epoch after the first 40 epochs. For each trajectory, we time-average the obtained samples and then average the results across trajectories. The SamAdams runs are executed first as the BAOAB stepsizes will be obtained from the adaptive stepsize values $\Delta t$ used by SamAdams (pooled from all trajectories). BAOAB is run with three different learning rates [^7] We use $h$ here to denote BAOAB’s stepsize to avoid confusion. $h$ (i.e., 300 BAOAB runs in total): The mean of the pooled $\Delta t$ (denoted by $h=\langle\Delta t\rangle$), the mean of the smallest 10% of the pooled $\Delta t$ (denoted by $h=\langle\Delta t\rangle_{\text{small}}$), and the mean of the largest 10% ($h=\langle\Delta t\rangle_{\text{large}}$). For reference, these values are $\langle\Delta t\rangle_{\text{small}}=0.00046$, $\langle\Delta t\rangle=0.00093$, and $\langle\Delta t\rangle_{\text{large}}=0.0014$. Note that this way of choosing the stepsizes of BAOAB is different from the last section where we took the point of view that a fair comparison between adaptive-stepsize and constant-stepsize schemes is made by considering the first 10% of obtained $\Delta t$ rather than the largest or smallest 10%. The results together with 95% confidence intervals are given in Table 2.
SamAdams BAOAB $h=\langle\Delta t\rangle_{\text{small}}$ BAOAB $h=\langle\Delta t\rangle$ BAOAB $h=\langle\Delta t\rangle_{\text{large}}$
Train Accuracy (%) 98.48$\pm$0.07 70.07$\pm$7.31 11.65$\pm$1.97 10.26$\pm$0.52
Test Accuracy (%) 97.97$\pm$0.05 70.26$\pm$7.33 11.63$\pm$1.98 10.23$\pm$0.53
Table 2: Mean accuracies and 95%-confidence intervals obtained by sampling the posterior of a CNN on the MNIST dataset. 100 independent trajectories were run per column. The three BAOAB stepsizes were obtained from the adaptive stepsize values used by SamAdams (see main text). The other hyperparameters are as in Fig. 17.
SamAdams reaches high accuracies with small variance, i.e. high reliability. Comparable to the single-trajectory results in Fig. 17, the two larger BAOAB stepsizes fail almost completely (since there are 10 classes, 10% corresponds to random class label assignment). Only when run at the smallest of the three chosen stepsizes does BAOAB train properly, but with far worse results than SamAdams. The fact that this stepsize only leads to ${}70\%$ accuracies when a similar stepsize led to accuracies of more than 95% in Fig. 17 implies that the variance at this stepsize must be high (also evident by the large confidence intervals in the table). In fact, when plotting the histograms of the (time-averaged) train accuracies of SamAdams and the $\langle\Delta t\rangle_{\text{small}}$-BAOAB runs, we see that while the SamAdams accuracies all lie north of 95%, BAOAB yields a significant number of accuracies close to 10%, see Fig. 18.
Figure 18: Histograms of time-averaged train accuracies for 100 independent trajectories of SamAdams (left) and BAOAB (right) corresponding to the results in Table 2. BAOAB was run at $h=\langle\Delta t\rangle_{\text{small}}=0.00046$, see main text for explanation.
6 Conclusion
We have presented a flexible integration framework for adaptive-stepsize Langevin sampling algorithms based on an auxiliary monitor variable. In particular, we have shown how the SamAdams algorithm, inspired by the Adam optimizer, shows superior behavior in terms of both stability and convergence speed compared to fixed stepsize alternatives. While we have provided various numerical experiments, we believe that there are many more settings in which the force-sensitive stepsize adaptation can greatly enhance sampling performance.
The method can be adapted to large-scale Bayesian machine learning, and is likely to show advantages in relation to models that are typically currently trained using Adam and its derivatives in the deep learning context [80], whether for natural language processing [78], diffusion models [51] or some other type of machine learning application. As sampling and BNN frameworks are introduced to address a wider range of AI challenges, we expect algorithms such as the one described here will be in high demand. Manual adjustment of the stepsize (learning rate scheduling) is often used in machine learning applications; the flexible nature of our framework suggests the possibility of an automated approach which can simplify training workflows if the right choice of monitor function and other aspects can be identified (which may not always be as simple as the norm of the gradient). The strong relation between the stepsize used in training, the batch size (amount of gradient noise) and the generalization error makes the new method interesting for active learning settings in which batch sizes vary in time [2]. Although it is not the main target of this work, we believe that SamAdams (or a similar method based on the adaptation framework presented in this article) might be of interest to computational scientists simulating physical models in which forces may increase preciptously during integration, often requiring the use of small stepsizes compared to the long simulation times that have to be realized.
Finally, we note that the SamAdams framework can easily be combined with other sampling procedures based on SDE discretization, since, as we have written it in Algorithm 1, the timestep adaptation is implemented separately from the propagation of state variables. It could also be combined with debiasing techniques (see [19]) to produce unbiased estimates from the target measure whilst still avoiding Metropolis-Hastings accept-reject steps.
Acknowledgements
This research was supported by the MAC-MIGS Centre for Doctoral Training (grant EPSRC EP/S023291/1). The authors wish to acknowledge Katerina Karoni for helpful discussions at an early stage of the project and Daniel Paulin for advice on the neural network studies. We also thank Gilles Vilmart for pointing out the article [99] which we have used to establish the weak convergence of our method and Michael Tretyakov for a helpful discussion on the numerical stability of SDE discretizations.
Appendix A General Form of the SamAdams Algorithm
It is possible to SamAdams-ize any integrator by wrapping it in Z-steps. For a standard form SDE
${\rm d}x_{t}=a(x_{t},t){\rm d}t+\sigma(x_{t},t){\rm d}W_{t},$ (26)
we can introduce auxiliary variable $\zeta$ controlled by
$\frac{{\rm d}{\zeta}_{\tau}}{{\rm d}\tau}=f(x_{\tau},\zeta_{\tau}),$
with, for example, $f(x_{\tau},\zeta_{\tau})=-\alpha\zeta_{\tau}+g(x_{\tau})$, then introduce a Sundman transformation
${\rm d}t=\psi(\zeta){\rm d}\tau,$
with $\psi$ a suitable uniformly positive, bounded, smooth function. Finally the rescaled equations corresponding to (26) become
${\rm d}x_{\tau}$ $=$ $\psi(\zeta)a(x_{\tau},t(\tau)){\rm d}{\tau}+\sigma(x_{\tau},t(\tau))\sqrt{\psi(\zeta)}{\rm d}W_{\tau},$ (27)
${\rm d}\zeta$ $=$ $f(x_{\tau},\zeta(\tau)){\rm d}\tau,$ (28)
${\rm d}t$ $=$ $\psi(\zeta){\rm d}\tau.$ (29)
[…]
$\zeta_{n+1/2}$ $=$ $\rho^{1/2}\zeta_{n}+\alpha^{-1}(1-\rho^{1/2})g(x_{n}),$ (30)
$\Delta t_{n+1}$ $=$ $\psi(\zeta_{n+1/2})\Delta\tau,$ (31)
$x_{n+1}$ $=$ $\hat{\Phi}_{\Delta t_{n+1}}(x_{n}),$ (32)
$\zeta_{n+1}$ $=$ $\rho^{1/2}\zeta_{n+1/2}+\alpha^{-1}(1-\rho^{1/2})g(x_{n+1}),$ (33)
$\mu_{n+1}$ $=$ $\psi(\zeta_{n+1}),$ (34)
with $\rho=\exp(-\alpha\Delta\tau)$ (compare Algorithm 1). In case $g$ depends on a force, that force calculation performed at the end of a step can be reused in the following step during the initial $Z$-half-step and in the state propagation. Like Alg. 1, this method uses the two half Z-steps to produce more accurate weights at the step endpoints.
Appendix B Adam Dynamics
In [22] and [50] it was demonstrated that the Adam optimizer can be interpreted as the Euler discretization of a certain system of ODEs, given by
$\mathop{}\!\text{d}x_{t}$ $=\frac{p_{t}}{\sqrt{\zeta_{t}+\epsilon}}\mathop{}\!\text{d}t,$ (35)
$\mathop{}\!\text{d}p_{t}$ $=-\nabla U(x_{t})\mathop{}\!\text{d}t-\gamma p_{t}\mathop{}\!\text{d}t,$ (36)
$\mathop{}\!\text{d}\zeta_{t}$ $=[\nabla U(x_{t})]^{2}\mathop{}\!\text{d}t-\alpha\zeta_{t}\mathop{}\!\text{d}t.$ (37)
Here we have $x_{t},p_{t},\zeta_{t}\in\mathbb{R}^{d}$, $\alpha>0$, $\gamma>0$, and the algebraic operations are to be understood elementwise. Applying an Euler discretization with step size ${\Delta t}$ leads to
$p_{n+1}$ $=(1-\gamma{\Delta t})p_{n}-{\Delta t}\nabla U(x_{n}),$ (38)
$\zeta_{n+1}$ $=(1-\alpha{\Delta t})\zeta_{n}+{\Delta t}[\nabla U(x_{n})]^{2},$ (39)
$x_{n+1}$ $=x_{n}+{\Delta t}\bigg{(}\frac{p_{n+1}}{\sqrt{\zeta_{n+1}+\epsilon}}\bigg{)}.$ (40)
Setting $\beta_{1}:=1-\gamma{\Delta t}\Rightarrow$ ${\Delta t}=(1-\beta_{1})/\gamma$ and $\beta_{2}:=1-\alpha{\Delta t}\Rightarrow{\Delta t}=(1-\beta_{2})/\alpha$, this becomes
$p_{n+1}$ $=\beta_{1}p_{n}-\frac{1-\beta_{1}}{\gamma}\nabla U(x_{n}),$ (41)
$\zeta_{n+1}$ $=\beta_{2}\zeta_{n}+\frac{1-\beta_{2}}{\alpha}[\nabla U(x_{n})]^{2},$ (42)
$x_{n+1}$ $=x_{n}+{\Delta t}\bigg{(}\frac{p_{n+1}}{\sqrt{\zeta_{n+1}+\epsilon}}\bigg{)}.$ (43)
Multiplying the $p$-equation by $\gamma$ and the $\zeta$-equation by $\alpha$, setting $\tilde{p}:=-\gamma p$ and $\tilde{\zeta}:=\alpha\zeta$, we obtain
$\tilde{p}_{n+1}$ $=\beta_{1}\tilde{p}_{n}+(1-\beta_{1})\nabla U(x_{n}),$ (44)
$\tilde{\zeta}_{n+1}$ $=\beta_{2}\tilde{\zeta}_{n}+(1-\beta_{2})[\nabla U(x_{n})]^{2},$ (45)
$x_{n+1}$ $=x_{n}-\widetilde{{\Delta t}}\bigg{(}\frac{\tilde{p}_{n+1}}{\sqrt{\tilde{\zeta}_{n+1}+\tilde{\epsilon}}}\bigg{)},$ (46)
where we also set $\widetilde{{\Delta t}}:=\frac{{\Delta t}\sqrt{\alpha}}{\gamma}$ and $\tilde{\epsilon}:=\alpha\epsilon$. Note that the new momentum variable is sign-flipped which leads to a plus sign in front of the force in the $\tilde{p}_{n+1}$-equation and a minus sign in front of the momentum in the $x_{n+1}$ equation, which is conventionally reversed for Langevin dynamics-based schemes. Apart from an additional scaling of $\tilde{p}_{n+1}$ and $\tilde{\zeta}_{n+1}$ (see the end of this section), (44)-(46) form the Adam scheme as introduced in [53] and employed by widely used machine learning frameworks such as PyTorch [73] or Tensorflow [1]. According to (45), the $n$-th iterate of $\tilde{\zeta}$ is given by
$\tilde{\zeta}_{n}=\beta_{2}^{n}\tilde{\zeta}_{0}+(1-\beta_{2})\sum_{i=0}^{n-1}\beta_{2}^{n-1-i}[\nabla U(x_{i})]^{2},$ (47)
which takes the form of an exponentially weighted average over the squared gradient components, where smaller weights are assigned to values further in the past. Since Adam is an optimizer, not a sampler, it is inherently deterministic. However, if one assumes the loss landscape itself is subject to noise (e.g., from dataset subsampling), such that in any given iteration, $U(x_{t})$ is an unbiased estimator of the true loss at that point, $\tilde{U}(x_{t})$, one observes that (47) estimates the second moments of the gradients, i.e., the uncentered variances [^8] This is only approximately true as the distribution of $U(x_{t})$ will be different for different points $x_{t}$ along the trajectory. However, as mentioned in Sec. 3 in [53], due to the decaying weights assigned to gradients further in the past in the sum in (47), the error can be assumed to be small.. Similarly, the momentum accumulates an estimate of the first moment, i.e., the expectation of the gradient. Without gradient noise, the moving averages $\tilde{p}_{n}$ and $\tilde{\zeta}_{n}$ can be interpreted as estimates of the averaged loss gradient and gradient variance. In both cases, the parameter update (46) approximates
$x_{n+1}=x_{n}-\widetilde{{\Delta t}}\bigg{(}\frac{\mathbb{E}\Big{[}\nabla U(x_{n})\Big{]}}{\sqrt{\text{Var}_{0}(\nabla U(x_{n}))+\tilde{\epsilon}}}\bigg{)},$ (48)
where the $\text{Var}_{0}(X)$ denotes the uncentered variance of $X$. One therefore obtains larger steps in regions of large gradients but small gradient variances (curvatures).
From (46) it can be seen that the adaptive stepsize in step $n$ given by
$\Delta t_{n}=\frac{\widetilde{{\Delta t}}}{\sqrt{\tilde{\zeta}_{n}+\tilde{\epsilon}}},$ (49)
such that larger values of the moving averages of the squared gradients lead to smaller adaptive stepsizes.
We note that Adam typically contains two additional steps: Before inserting the momentum $\tilde{p}_{n+1}$ and stepsize-scaling variable $\tilde{\zeta}_{n+1}$ into (46), they are rescaled according to
$\hat{p}_{n+1}=\frac{\tilde{p}_{n+1}}{1-\beta_{1}^{n+1}},\hskip 28.45274pt\hat{\zeta}_{n+1}=\frac{\tilde{\zeta}_{n+1}}{1-\beta_{2}^{n+1}}.$ (50)
This is done in order to remove the bias due to the initial conditions from the estimates of the gradient moments (see Algorithm 1 and Sec. 3 in [53]). Since both $\beta_{1}^{n}\to 0$ and $\beta_{2}^{n}\to 0$ for $n\to\infty$, these steps can often be skipped in practice, which is why we will not consider them in the rest of this work.
Appendix C Further Details of Stepsize Control Mechanism
The stepsize adaptation mechanism used by the SamAdams process presented in Sec. 4 in the main text is based on three components. The first is the evolution of the force-sensitive $\zeta$-variable, given by
$\mathop{}\!\text{d}\zeta_{\tau}=-\alpha\zeta_{\tau}\mathop{}\!\text{d}\tau+\Omega^{-1}\|\nabla U(x_{\tau})\|^{s}\mathop{}\!\text{d}\tau,$ (51)
with the two hyperparameters $\alpha>0$ and $\Omega>0$, whose role will be expanded upon below. The second component is given by the choice of a Sundman transform kernel $\psi$ as a function of $\zeta$. For example, we set $\psi(\zeta)\equiv\psi^{(1)}(\zeta)$ with $\psi^{(1)}$ from Sec. 4, given by
$\psi^{(1)}(\zeta)=m\frac{\zeta^{r}+M}{\zeta^{r}+m},\hskip 10.0pt\text{with }\hskip 2.0pt0<m<M<\infty.$ (52)
As mentioned in the main text, this form resembles the term used by Adam but allows for more flexibility via the scaling hyperparameter $r>0$ and stability due to its boundedness, $\psi^{(1)}(\zeta)\in(m,M)$ for all $\zeta>0$ (note that once $\zeta$ is initialized to $\zeta_{0}>0$, it will always remain positive due to (51)). The third component is the time rescaling relationship $\Delta t=\psi(\zeta)\Delta\tau$, which scales the constant stepsize in virtual time, $\Delta\tau$, with the transform kernel evaluated at $\zeta$ to yield the adaptive stepsize in real time, $\Delta t$. Fig. 19 shows the transform kernel $\psi^{(1)}(\zeta)$ with $m=0.1$ and $M=10$ (the values that were used in most of our experiments) and for various $r$. Since $m$ and $M$ are the bounds of $\psi^{(1)}$, they also specify the bounds on the adaptive stepsize, $\Delta t\in(m\Delta\tau,M\Delta\tau)$. As denoted by the black arrows in the figure, larger forces tend to increase the value of $\zeta$ which in turn decreases the value of $\psi^{(1)}(\zeta)$, leading to a decrease in adaptive stepsize. The red line denotes the value of $\zeta$ for $\psi^{(1)}(\zeta)=1$, i.e., where $\Delta t=\Delta\tau$. It thus gives the boundary between the $\zeta$-regions where $\Delta t$ is smaller or greater than $\Delta\tau$. For $m=1/M$, this boundary is exactly at $\zeta=1$, i.e., $\psi^{(1)}(\zeta)=1$ (the red dot). Note that this changes for different $m$ and $M$. The exponent $r$ governs the sensitivity of the overall mechanism.
Figure 19: Time transform kernel $\psi(\zeta)\equiv\psi^{(1)}(\zeta)$ as a function of $\zeta$ for $m=0.1$, $M=10$, and different exponents $r$. The black dashed lines denote the bounds $m$ and $M$. The red point at (1,1) and red dashed line separate the $\zeta$-region in which the basic stepsize $\Delta\tau$ is magnified from the region where it is reduced (i.e., the regions of $\psi^{(1)}(\zeta)>1$ and $\psi^{(1)}(\zeta)<1$, resp.). The arrows denote the direction of the $\zeta$-evolution dependent on the experienced forces $\|\nabla U\|^{s}$.
How large the forces need to be to lead to a decrease of $\Delta t$ below $\Delta\tau$ (for a fixed set of transform kernel parameters $r,m,M$) can be controlled with the hyperparameters $\alpha$ and $\Omega$ in (51). From looking at the solution of (51) given by (see(16) in the main text)
$\zeta(\tau)=e^{-\alpha\tau}\zeta(0)+\Omega^{-1}\int_{0}^{\tau}e^{-\alpha(\tau-s)}\|\nabla U(x_{s})\|^{s}{\rm d}s,$ (53)
we see that that $\zeta$ is identical to an exponentially weighted moving average over the past force magnitudes raised to the power $s$, where $\alpha$ governs how strongly past values are suppressed. Pulling $\Omega^{-1}$ into the integral, it is clear that we actually average over $\Omega^{-1}\|\nabla U\|^{s}$, which means that $\Omega^{-1}$ linearly scales the obtained $\zeta$ values, directly influencing the size of $\Delta t$. For a more concrete view on the influence of the two parameters, we look at the discretized version of (53) employed by our splitting integrators described in Sec. 3 in the main text, given by
$\zeta_{n+1}=\Phi_{\Delta\tau}^{{\rm Z}}(\zeta_{n},x_{n})=e^{-\Delta\tau\alpha}\zeta_{n}+\frac{1}{\Omega\alpha}(1-e^{-\Delta\tau\alpha})\|\nabla U(x_{n})\|^{s}.$ (54)
From this, it follows that the $n$-th iterate is given by
$\zeta_{n}=p^{n}\zeta_{0}+q\bigg{(}\sum_{i=1}^{n}p^{n-i}\|\nabla U(x_{i})\|^{s}\bigg{)},$ (55)
with $p:=\exp(-\Delta\tau\alpha)$, $q:=\frac{1}{\Omega\alpha}(1-p)$. This form is equivalent to the one used in Adam, see (47). As mentioned in the main text, the fundamental difference is that in Adam the $\tilde{\zeta}$ in (47) is a vector and the squaring of the force is an elementwise operation (leading to an adaptive stepsize per degree of freedom), which in our case is replaced by the Euclidean norm of the force, leading to a scalar $\zeta$ (and hence a single adaptive stepsize for all degrees of freedom). We confirm again that $\alpha$ through $p$ governs the influence of past forces, suppressing the ones further in the past more strongly. Larger values for $\alpha$ lead to smaller weights assigned to past forces leading to less "memory" in the $\zeta$-dynamics. The parameter $q$ linearly scales the contribution of the sum and thus the overall magnitude of $\zeta$ (and thus the obtained adaptive stepsize $\Delta t$). Note that $q$ depends on both $\alpha$ and $\Omega$, which means varying $\alpha$ will not only change the influence of past force values but also change the overall magnitude of $\zeta$ If we simply wanted to change the moving average behavior without changing the average $\Delta t$-level, we have to adjust $\alpha$ such that $\mathbb{E}(\zeta_{n})$ is kept constant, where the expectation is with respect to the (unknown) evolving law of $\zeta_{n}$. We care only about what happens at long times. Assume we have $(x_{0},p_{0},\zeta_{0})$ sampled from the (unknown) invariant measure of the SamAdams dynamics. Since our proposed integrator (Algorithm 1 in the main text) approximates the invariant measure, we may assume $\mathbb{E}(\zeta_{n})\approx E_{\zeta}$ and $\mathbb{E}\Big{[}\|\nabla U(x_{n})\|^{s}\Big{]}\approx E_{U}$ for all $n$. Taking the expectation of (55), we then have
$E_{\zeta}\approx p^{n}E_{\zeta}+q\bigg{(}\sum_{i=1}^{n}p^{n-i}E_{U}\bigg{)}.$ (56)
Using properties of the geometric series and the definition of $q$, we obtain
$E_{\zeta}\approx\frac{1}{\Omega\alpha}E_{U}.$ (57)
Thus, if we change $\alpha$ to influence the amount of memory in the system but we want to preserve the mean adaptive stepsize $\mathbb{E}(\Delta t)$, we need to set
$\Omega^{\text{new}}=\frac{\alpha^{\text{old}}}{\alpha^{\text{new}}}\Omega^{\text{old}},$ (58)
keeping the product $\Omega\alpha$ constant. Meanwhile, if we simply want to change the average magnitude of $\zeta$ (and hence $\Delta t$) without changing the weights of the moving average, we simply change $\Omega$ while keeping $\alpha$ fixed. In general terms, we obtain the following rules of thumb.
  1. 1.
    For fixed $\alpha$, smaller values of $\Omega$ lead to larger $\zeta$ and hence smaller $\Delta t$.
  2. 2.
    For fixed $\Omega$, larger values of $\alpha$ lead to smaller $\zeta$ and hence larger $\Delta t$, and also to less memory in $\zeta$.
  3. 3.
    Larger values of $\alpha$ with $\Omega$ scaled according to (58) leads to less memory in $\zeta$ while keeping $\mathbb{E}(\Delta t)$ fixed.
In Fig. 20, we demonstrate the effects of changing $\alpha$ and $\Omega$ on the obtained stepsizes on the 2-dimensional star potential (see Fig. 1 in the main text).
Figure 20: Adaptive stepsize obtained by SamAdams on a 2D test problem. Left: Changing $\alpha$ changes the smoothness of $\Delta t$. If $\Omega\alpha$ is identical for two different values of $\alpha$, $\mathbb{E}(\Delta t)$ is roughly identical as well. Right: Changing $\Omega$ while keeping $\alpha$ fixed changes $\mathbb{E}(\Delta t)$ without influencing its smoothness. The red dashed lines denote the minimum admissible stepsize, i.e., $\Delta t=m\Delta\tau$ (here $\Delta\tau=0.01,\ m=10$.)
In the left-hand plot, we start with the orange curve obtained by $\alpha_{1}=0.001$, $\Omega=100$, and hence $\Omega\alpha=0.1$. Increasing the value of $\alpha$ by a factor of 10 and simultaneously changing $\Omega^{-1}$ by the same factor, such that $\Omega\alpha$ remains the same, we decrease the memory in the $\zeta$-dynamics leading to more spontaneous changes in $\Delta t$ while keeping the overall magnitude of $\Delta t$ fixed (blue curve). Increasing $\alpha$ without decreasing $\Omega$ will also remove memory from the system but in a less controlled way, as one also obtains different $\mathbb{E}(\Delta t)$ (purple curve). On the right-hand side, starting again with the orange curve, we see that an increase of $\Omega$ while keeping $\alpha$ fixed leads to a systematic increase of $\Delta t$ without changing the moving average properties of the system. These mechanisms also work on more complex examples such as neural networks, see Fig. 21.
Figure 21: Adaptive stepsize trajectories of SamAdams for a simple fully connected neural network (one hidden layer) on a spiral/Swiss-roll-type binary classification task. Top: Effect of changing $\alpha$ on $\Delta t$ if $\Omega\alpha$ remains fixed. Bottom: Effect of changing $\Omega$ if $\alpha$ remains fixed. Hyperparameters: $T=\gamma=1$, $\Delta\tau=0.03$, $m=0.1$, $M=10$, $r=0.25$, $s=2$.
Note that the rules (1)-(3) above can only be regarded as rules of thumb. They do not strictly hold near the stability threshold (close to which anything can happen). Even for stepsizes well below the threshold, the rules only hold for changes in $(\alpha,\Omega)$ that don’t lead to qualitatively different trajectories. For example, increasing $\Omega$ while keeping $\alpha$ fixed will tend to increase $\Delta t$. A trajectory with higher $\Delta t$ might then experience larger forces compared to before, which will then lead to a decrease of $\Delta t$ again by virtue of larger obtained $\zeta$-values.
We further remark that on data science problems where the force $\nabla U$ is given by a sum over the data points, we find that $\Omega$ should often be chosen as $\mathcal{O}(N)$ or $\mathcal{O}(N^{s})$ with $N$ the number of data points. This prevents $\zeta$ from drifting off to too large values which would lead to $\Delta t_{n}=m\Delta\tau\ \text{for all }n.$ Note that the loss gradient is often normalized by the number of data points by default in common machine learning packages, which might render this point obsolete.
Finally, it is also instructive to consider the limiting cases of (54). We have:
$\lim_{\alpha\to\infty}\zeta_{n+1}=0,$ (59)
which leads to $\Delta t_{n+1}=M\Delta\tau$ for all $n$ (maximum $\Delta t$),
$\lim_{\alpha\to 0}\zeta_{n+1}=\zeta_{n}+\Omega^{-1}\Delta\tau\|\nabla U(x_{n})\|^{s},$ (60)
such that $\zeta_{n}\to\infty$ for $n\to\infty$ and hence $\Delta t_{n}\to m\Delta\tau$ (minimum $\Delta t$),
$\lim_{\Omega\to 0}\zeta_{n+1}=\infty,$ (61)
and hence $\Delta t_{n+1}=m\Delta\tau$ for all $n$ (minimum $\Delta t$), and
$\lim_{\Omega\to\infty}\zeta_{n+1}=e^{-\Delta\tau\alpha}\zeta_{n},$ (62)
leading to $\zeta_{n}\to 0$ for $n\to\infty$ and hence $\Delta t_{n}\to M\Delta\tau$.
An interesting case arises for the limit $\alpha\to\infty$ such that $\Omega\alpha$ is kept constant,
$\lim_{\begin{subarray}{c}\alpha\to\infty,\\ \Omega\alpha=\text{const}\end{subarray}}\zeta_{n+1}=\frac{1}{\Omega\alpha}\|\nabla U(x_{n})\|^{s}.$ (63)
This is a case with no memory but finite $\zeta$-values, where the adaptive stepsize is obtained via
$\Delta t=\psi(\|\nabla U(x)\|^{sr})\Delta\tau,$ (64)
which corresponds to the method used in [61].
Appendix D Example with Gradient Noise in Logistic Regression
We sample from the posterior of a logistic regression model on the forest covertype dataset [8]. This time, we examine the effect of using stochastic gradients with various batch sizes. We run SamAdams with various batch size and measure the resulting training accuracies. To compare it with BAOAB, we measure the mean adaptive stepsize of SamAdams on the full-batch run and execute all BAOAB runs at this fixed stepsize. This setting captures how the two samplers cope with changing batch sizes (leaving all other hyperparameters fixed). Fig. 22 shows the results for 4 different values of $\alpha$.
Figure 22: Training accuracies against stochastic gradient batch size $B$ for 4 different values of SamAdams hyperparameter $\alpha$. The stepsize of BAOAB was chosen to be identical to the mean of the adaptive stepsizses adopted by SamAdams on the full batch run on the given $\alpha$.
As the batch size decreases, the accuracy in both methods drops. However, due to the gradient noise, the otherwise simple dynamics (logistic regression posterior is convex) becomes more unstable, leading to automatic stepsize reduction in SamAdams. The sampling bias introduced by gradient noise from data subsampling scales with the stepsize, so a reduction of stepsize for decreasing batch size can restore accuracy. Note that the performance of BAOAB can be restored by picking a smaller stepsize as well. We merely wish to highlight the benefit of SamAdams to automatically react to induced gradient noise. Fig. 23 shows the batch size-dependent $\Delta t$-histograms obtained by SamAdams.
Figure 23: SamAdams (ZBAOABZ) $\Delta t$ histograms against batch size for the experiment of Fig. 22.
We see a decrease of the average stepsize with decreasing batch size, consistent across all $\alpha$ values. The mean and variance (and even the type of the distribution) depends on the hyperparameters $\alpha$ and $\Omega$ (with $\Omega$ kept constant here). To illustrate the sensitivity of the stepsize adaptation on gradient noise, we dynamically switch the batch size between 1 and 1,000 samples and inspect the behavior of the stepsize $\Delta t$. Fig. 24 shows the result, together with curves for constant batch size $B=1$ and $B=1,000$.
Figure 24: Adaptive stepsize of SamAdams when run with different batches. The green curve uses alternating batch sizes with $B$ changing between 1 and 1,000. The curves were averaged over 100 independent trajectories.
The stepsize adaptation is able to change $\Delta t$ dynamically and almost instantaneously in accordance with the batch size. While we did not find significant differences in the resulting accuracies of SamAdams and BAOAB when using dynamically changing batch sizes, we believe that this is a consequence of the simplicity of the problem. Logistic regression on comparably small datasets does not require stochastic gradients, neither for efficiency nor to escape local minima. It would be interesting to examine the combined impact of stepsize adaptation and gradient noise on large-scale deep learning models, where gradient subsampling is quintessential for efficient training, or in active learning settings, in which the batch size of different training iterations is allowed to vary.
Appendix E CNN Architecture
The following code specifies the simple convolutional neural network used for the MNIST experiments in Sec. 5.3.4 in the main text, implemented in PyTorch.
class SimpleCNN(nn.Module):
def __init__(self):
super(SimpleCNN, self).__init__()
self.conv1 = nn.Conv2d(in_channels=1, out_channels=32, kernel_size=3, padding=1)
self.pool = nn.MaxPool2d(kernel_size=2, stride=2)
self.conv2 = nn.Conv2d(in_channels=32, out_channels=64, kernel_size=3, padding=1)
self.conv3 = nn.Conv2d(in_channels=64, out_channels=128, kernel_size=3, padding=1)
self.fc_input_size = 128 * 3 * 3 # based on MNIST image dimension.
self.fc1 = nn.Linear(self.fc_input_size, 512)
self.fc2 = nn.Linear(512, 256)
self.fc3 = nn.Linear(256, 10)
def forward(self, x):
x = self.pool(nn.ReLU()(self.conv1(x)))
x = self.pool(nn.ReLU()(self.conv2(x)))
x = self.pool(nn.ReLU()(self.conv3(x)))
x = x.view(x.size(0), -1)
x = nn.ReLU()(self.fc1(x))
x = nn.ReLU()(self.fc2(x))
x = self.fc3(x)
x = torch.log_softmax(x, dim=1)
return x